### v1

# Neural Flow Diffusion Models: Learnable Forward Process for Improved Diffusion Modelling

Abstract
Conventional diffusion models typically relies on a fixed forward process, which implicitly defines complex marginal distributions over latent variables. This can often complicate the reverse process’ task in learning generative trajectories, and results in costly inference for diffusion models. To address these limitations, we introduce Neural Flow Diffusion Models (NFDM), a novel framework that enhances diffusion models by supporting a broader range of forward processes beyond the fixed linear Gaussian. We also propose a novel parameterization technique for learning the forward process. Our framework provides an end-to-end, simulation-free optimization objective, effectively minimizing a variational upper bound on the negative log-likelihood. Experimental results demonstrate NFDM’s strong performance, evidenced by state-of-the-art likelihood estimation. Furthermore, we investigate NFDM’s capacity for learning generative dynamics with specific characteristics, such as deterministic straight lines trajectories. This exploration underscores NFDM’s versatility and its potential for a wide range of applications.
diffusion, generative models, variational inference
1 Introduction
Generative models are a versatile class of probabilistic machine learning models, which finds applications across diverse fields such as art, music, medicine, and physics (Tomczak, 2022; Creswell et al., 2018; Papamakarios et al., 2021; Yang et al., 2022). Generative models excel in constructing probability distributions that accurately represent datasets, enabling them to produce new samples akin to the original data. Such capabilities make them ideal for tasks like enhancing training data through the generation of synthetic datasets (data augmentation) and uncovering hidden patterns and structures in an unsupervised learning setting.
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of generative models are constructed by two key processes: the forward process and the reverse process. The forward process gradually corrupts the data distribution, transforming it from its original form to a noised state. The reverse process learns to invert corruptions of the forward process and restore the data distribution. This way, it learns to generate data from pure noise. Diffusion models have demonstrated remarkable results in various domains (Ho et al., 2022; Saharia et al., 2022; Tan et al., 2024; Watson et al., 2023; Trippe et al., 2023). Nevertheless, most existing diffusion models fix the forward process to be predefined Gaussian, which makes it unable to adapt to the task at hand or simplify the target for the reverse process. At the same time there are many works that demonstrate how modifications of the forward process improve performance in terms of generation quality (Nichol & Dhariwal, 2021; Vahdat et al., 2021; Daras et al., 2022), likelihood estimation (Kingma et al., 2021; Nielsen et al., 2023; Bartosh et al., 2023) or sampling speed (Lee et al., 2023; Pooladian et al., 2023; Tong et al., 2023).
In this paper, we present Neural Flow Diffusion Models (NFDM), a framework that allows for the pre-specification and learning of latent variable distributions defined by the forward process. Unlike conventional diffusion models (Ho et al., 2020), which rely on a conditional Gaussian forward process, NFDM may accommodate any continuous (and learnable) distribution that can be expressed as an invertible mapping applied to noise. We also derive, and leverage, an end-to-end simulation-free optimization procedure, that minimizes a variational upper bound on the negative log-likelihood (NLL).
We also propose an efficient neural network-based parameterization for the forward process, enabling it to adapt to the reverse process during training and simplify the learning of the data distribution. To demonstrate NFDM’s capabilities with a learnable forward process we provide experimental results on CIFAR-10, ImageNet 32 and 64, attaining state-of-the-art NLL results, which is crucial for many applications such as data compression (Ho et al., 2021; Yang & Mandt, 2022), anomaly detection (Chen et al., 2018b; Dias et al., 2020) and out-of-distribution detection (Serrà et al., 2019; Xiao et al., 2020).
Leveraging the flexibility of NFDM, we further explore training with constraints on the reverse process to learn generative dynamics with specific properties. As a case study, we discuss a curvature penalty on the deterministic generative trajectories. Our empirical results indicate improved computational efficiency compared to baselines on synthetic datasets as well as MNIST, CIFAR-10, and downsampled ImageNet.
We summarize our contributions as follows:
  1. 1.
    We introduce Neural Flow Diffusion Models (NFDM), improving diffusion modelling through a learnable forward process.
  2. 2.
    We develope an end-to-end optimization procedure that minimizes an upper bound on the negative log-likelihood in a simulation-free manner.
  3. 3.
    We demonstrate state-of-the-art log-likelihood results on CIFAR-10, ImageNet 32 and 64.
  4. 4.
    We show how NFDM can be used in learning generative processes with specific properties, exemplified by dynamics with straight line trajectories, where NFDM leads to significantly faster sampling speeds and enhanced generation quality with fewer sampling steps.
2 Background
Diffusion models are generative latent variable models consisting of two processes: the forward and the reverse (or generative) process. The forward process is a dynamic process that takes a data point $\mathbf{x}\sim q(\mathbf{x}),\mathbf{x}\in\mathbb{R}^{D}$, and perturbs it over time by injecting noise. This generates a trajectory of latent variables $\{\mathbf{z}(t)\}_{t\in[0,1]}$, conditional on the data $\mathbf{x}$, where $[0,1]$ is a fixed time horizon and $\mathbf{z}_{t}=\mathbf{z}(t)\in\mathbb{R}^{D}$. The (conditional) distribution can be described by an initial distribution $q(\mathbf{z}_{0}|\mathbf{x})$ and a Stochastic Differential Equation (SDE) with a linear drift term $\tilde{f}^{F}(\mathbf{z}_{t},t):\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$, scalar variance $g(t):[0,1]\mapsto\mathbb{R}_{+}$, and a standard Wiener process $\mathbf{w}$:
$d\mathbf{z}_{t}$ $=\tilde{f}^{F}(\mathbf{z}_{t},t)dt+g(t)d\mathbf{w}.$ (1)
Due to the linearity of $\tilde{f}^{F}$, we can reconstruct the conditional marginal distribution $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}I)$. Typically, the conditional distributions evolve from some low variance distribution $q(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{z}_{0}-\mathbf{x})$ to a unit Gaussian $q(\mathbf{z}_{1}|\mathbf{x})\approx\mathcal{N}(\mathbf{z}_{1};0,I)$. This forward process is then reversed by starting from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and following the reverse SDE (Anderson, 1982):
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}(\mathbf{z}_{t},t)dt+g(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (2)
$\tilde{f}^{B}(\mathbf{z}_{t},t)$ $=\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}).$ (3)
Here, $\bar{\mathbf{w}}$ denotes a standard Wiener process where time flows backwards. Diffusion models approximate this reverse process by learning $\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t})$, known as the score function, through a $\lambda_{t}$-weighted score matching loss:
$\underset{u(t)q(\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t})\big{\|}_{2}^{2}\right],$ (4)
where $u(t)$ represents a uniform distribution over the interval $[0,1]$, and $s_{\theta}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ is a learnable approximation. This loss can be reformulated as a more tractable denoising score matching loss Vincent (2011):
$\underset{u(t)q(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\big{\|}_{2}^{2}\right].$ (5)
With a learned score function $s_{\theta}(\mathbf{z}_{t},t)$, one can generate a sample from the reverse process by first sampling from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and then simulating the reverse SDE:
$d\mathbf{z}_{t}$ $=\big{[}\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)s_{\theta}(\mathbf{z}_{t},t)\big{]}dt+g(t)d\bar{\mathbf{w}},$ (6)
resulting a sample $\mathbf{z}_{0}\sim p_{\theta}(\mathbf{z}_{0})\approx q(\mathbf{z}_{0})\approx q(\mathbf{x})$.
Diffusion models possess several important properties. For example, for specific $\lambda_{t}$, the objective (eq. 5) can be reformulated (Song et al., 2021) as an Evidence Lower Bound (ELBO) on the model’s likelihood. Furthermore, the minimization of denoising score matching (eq. 5) is a simulation-free procedure. This means that simulating either the forward or reverse processes through its SDE is not necessary for sampling $\mathbf{z}_{t}$, nor is it necessary for estimating the gradient of the loss function. Instead, we can directly sample $\mathbf{z}_{t}\sim q(\mathbf{z}_{t}|\mathbf{x})$. The simulation-free nature of this approach is a crucial aspect for efficient optimization.
Another notable property is the existence of an Ordinary Differential Equation (ODE) corresponding to the same marginal densities $q(\mathbf{z}_{t})$ as the SDE (eq. 1):
$d\mathbf{z}_{t}$ $=f(\mathbf{z}_{t},t)dt,\quad\textrm{where}$ (7)
$f(\mathbf{z}_{t},t)$ $=\tilde{f}^{F}(\mathbf{z}_{t},t)-\frac{g^{2}(t)}{2}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}).$ (8)
This implies that we can sample from diffusion models deterministically, allowing the use of off-the-shelf numerical ODE solvers for sampling, which may improve the sampling speed compared to stochastic sampling that requires simulating an SDE. Additionally, deterministic sampling enables us to compute densities by treating the model as a continuous normalizing flow, as detailed in Chen et al. (2018a); Grathwohl et al. (2018).
3 Neural Flow Diffusion Models
Diffusion models can be viewed as a specific type of hierarchical Variational Autoencoders (VAEs) (Kingma & Welling, 2013; Rezende et al., 2014), where the latent variables are inferred through a pre-specified scaling of the data point and injection of Gaussian noise. However, this formulation limits diffusion models in terms of the flexibility of their latent space, and makes learning of the reverse process more challenging. To address this limitation, we propose a generalized form of data and noise transformations that enables the definition and learning of a broad range of distributions in the latent space beyond the conventional Gaussian.
In this section, we introduce Neural Flow Diffusion Models (NFDM)– a framework that generalizes conventional diffusion models. The key idea in NFDM is to define the forward process implicitly via a learnable transformation $F_{\varphi}(\varepsilon,t,\mathbf{x})$. This lets the user define a broad range of continuous time and data dependent forward process. Importantly, NFDM retains crucial properties of conventional diffusion models, like likelihood-based and simulation-free training. Previous diffusion models emerge as special cases when the data transformation is linear, time-independent, and/or additive Gaussian. The connections between NFDM and existing models are discussed in Section 6.
3.1 Forward process
The forward process defines a target for the reverse process. It describes the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ in the latent space that the reverse process tries to match.
We approach the forward process constructively. First, we define a conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for $t\in[0,1]$. Then, we introduce the corresponding conditional ODE that together with the initial distribution $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ matches the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. Finally, we define a conditional SDE that implicitly defines distribution over trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, with the same marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
Forward Marginal Distribution. We characterize the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ of the forward process trough a function of injected noise $\varepsilon$, time $t$, and the data $\mathbf{x}$:
$\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x}),$ (9)
where $F_{\varphi}:\mathbb{R}^{D}\times[0,1]\times\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$ and $\varepsilon\sim q(\varepsilon)=\mathcal{N}(\varepsilon;0,I)$. This function transforms a sample from a distribution $q(\varepsilon)$ into $\mathbf{z}_{t}$, conditional on the data point $\mathbf{x}$ and time step $t$, implicitly defining the conditional distribution of latent variables $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$. Additionally, eq. 9 facilitates direct and efficient sampling from $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ through $F_{\varphi}$.
Conditional ODE. Assuming that $F_{\varphi}$ is differentiable with respect to $\varepsilon$ and $t$ and invertible with respect to $\varepsilon$, fixing specific values of $\mathbf{x}$ and $\varepsilon$ and varying $t$ from 00 to $1$ results in a smooth trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}_{1}$. Differentiating these trajectories over time yields a velocity field corresponding to the conditional distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, thereby defining a conditional ODE:
$d\mathbf{z}_{t}$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt,\quad\textrm{where}$ (10)
$f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=\left.\frac{\partial F_{\varphi}(\varepsilon,t,\mathbf{x})}{\partial t}\right|_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}.$ (11)
Therefore, if we sample $\mathbf{z}_{0}\sim q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ and solve the ODE in eq. 10 until time $t$, we have a sample $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
The time derivative of $F_{\varphi}$ may be calculated using automatic differentiation tools like PyTorch (Paszke et al., 2017) or JAX (Bradbury et al., 2018). Specifically, we can compute a Jacobian Vector Product (JVP) with a unit one-dimensional vector. Unlike standard backpropagation, which corresponds to a Vector Jacobian Product (VJP), JVP does not necessitate $\mathcal{O}(D)$ backpropagation passes, enhancing scalability.
Conditional SDE. The function $F_{\varphi}$ and the distribution of the noise $q(\varepsilon)$ together defines $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$. To completely define the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, we introduce a conditional SDE that starts from sample $\mathbf{z}_{0}$ and runs forward in time.
With access to both the ODE and score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, the SDE (Song et al., 2020b) corresponding to the distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, is:
$d\mathbf{z}_{t}$ $=\tilde{f}^{F}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w},\quad\textrm{where}$ (12)
$\tilde{f}^{F}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})+\frac{g_{\varphi}^{2}(t)}{2}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x}).$
Here, $g_{\varphi}:[0,1]\mapsto\mathbb{R}_{+}$ is a scalar function, and $\mathbf{w}$ represents a standard Wiener process. It is important to note, that $g_{\varphi}$ only influences the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, not the marginal distributions. The Conditional SDE in eq. 12 has the same conditional marginal distribution, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, for any $g_{\varphi}$.
The score function of $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is:
$\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\nabla_{\mathbf{z}_{t}}\left[\log q(\varepsilon)+\log\left|J_{F}^{-1}\right|\right]$ (13)
$\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})\quad\textrm{and}\quad J_{F}^{-1}=\frac{\partial F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})}{\partial\mathbf{z}_{t}}.$ (14)
While the log-determinant of the Jacobian matrix $J_{F}^{-1}$ is not generally available in explicit form, it can be analytically evaluated for certain transformations or specific architectures of the function $F_{\varphi}$, like RealNVP style architectures (Dinh et al., 2016; Kingma & Dhariwal, 2018). The parameterization of $F_{\varphi}$ is further discussed in Section 3.5.
Given access to the log probability $\log q(\varepsilon)$ and the log-determinant of $J_{F}^{-1}$, automatic differentiation tools enable the calculation of gradients and the score function. Unlike the time derivative, regular backpropagation can be utilized here, ensuring computationally efficient score function evaluation.
3.2 Reverse process
To define the reverse (generative) process, we specify a reverse SDE that starts from $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$ and runs backwards in time. To do so we first introduce a conditional reverse SDE that reverses the conditional forward SDE (eq. 12). Following (Anderson, 1982), we define:
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w},\quad\textrm{where}$ (15)
$\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\frac{g_{\varphi}^{2}(t)}{2}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x}).$
Secondly, we reference the reverse SDE (see eq. 15) and incorporating the prediction of $\mathbf{x}$:
$d\mathbf{z}_{t}=$ $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (16)
$\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)},$ (17)
and $\hat{\mathbf{x}}_{\theta}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ is a function that predicts the data point $\mathbf{x}$. This SDE defines the dynamics of the generative trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$.
However, to fully specify the reverse process, it is necessary to define a prior distribution $p(\mathbf{z}_{1})$ and a reconstruction distribution $p(\mathbf{x}|\mathbf{z}_{0})$. In all of our experiments, we set the prior $p(\mathbf{z}_{1})$ to be a unit Gaussian distribution $\mathcal{N}(\mathbf{z}_{1};0,I)$ and let $p(\mathbf{x}|\mathbf{z}_{0})$ be a Gaussian distribution with a small variance $\mathcal{N}(\mathbf{x};\mathbf{z}_{0},\delta^{2}I)$, where $\delta^{2}=10^{-4}$.
This particular parameterization of the reverse process is not the only possibility. However, it is a convenient choice as it allows for the definition of the reverse process simply through the prediction of $\mathbf{x}$, akin to conventional diffusion models (Ho et al., 2020).
3.3 Optimization
We propose to optimize the forward and reverse processes of NFDM jointly, minimizing the following objective:
$\mathcal{L}={\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}},$ (18)
${\mathcal{L}_{\mathrm{rec}}}=\underset{q_{\varphi}(\mathbf{x},\mathbf{z}_{0})}{\mathbb{E}}\left[-\log p(\mathbf{x}|\mathbf{z}_{0})\right],$ (19)
${\mathcal{L}_{\mathrm{diff}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right],$ (20)
${\mathcal{L}_{\mathrm{prior}}}=\underset{q(\mathbf{x})}{\mathbb{E}}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})\|p(\mathbf{z}_{1})\big{)}\right].$ (21)
As we demonstrate in Section A.1 this objectives provide a variational bound on the model’s likelihood $p_{\theta,\varphi}(\mathbf{x})$.
The objective $\mathcal{L}$ shares similarities with the standard diffusion model objective (Ho et al., 2020), though the individual terms differ. Moreover, if the forward process is parameterized by $\varphi$, we need to optimize the objective (eq. 18) with respect to these parameters. The objective $\mathcal{L}$ also exhibits strong connections with both the Flow Matching (Lipman et al., 2022) and Score Matching (Vincent, 2011) objectives. We explore these connections in Section B.2, where we also discuss the role of $g_{\varphi}$.
A key characteristic of the NFDM objective is its amenability to training within a simulation-free paradigm, which is critical for efficient optimization. We summarize the training procedure in Algorithm 1.
Algorithm 1 Optimization of NFDM
0:  $q(x)$, $F_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$
  for learning iterations do
     $\mathbf{x}\sim q(x)$, $t\sim u(t)$, $\{\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{1}\}\sim q(\cdot|\mathbf{x})$
     ${\mathcal{L}_{\mathrm{rec}}}=-\log p(\mathbf{x}|\mathbf{z}_{0})$
     ${\mathcal{L}_{\mathrm{diff}}}=\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}$
     ${\mathcal{L}_{\mathrm{prior}}}=\log q_{\varphi}(\mathbf{z}_{1}|x)-\log p(\mathbf{z}_{1})$
     Gradient step on $\theta$ and $\varphi$ w.r.t. ${\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}}$
  end for
Algorithm 2 Stochastic sampling from NFDM
0:  $F_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$, $T$ – number of steps
  $\Delta t=\frac{1}{T}$, $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$
  for $t=1,\dots,\frac{2}{T},\frac{1}{T}$ do
     $\bar{\mathbf{w}}\sim\mathcal{N}(0,I)$
     $\mathbf{z}_{t-\Delta t}=\mathbf{z}_{t}-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\Delta t+g_{\varphi}(t)\bar{\mathbf{w}}\sqrt{\Delta t}$
  end for
  $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$
3.4 Sampling
NFDM offers several methods for sampling from the trained reverse process. The first method is stochastic sampling, as defined in Section 3.2. This process involves the following steps: 1) sample $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$; 2) numerically simulate the SDE (eq. 16) to obtain $\mathbf{z}_{0}$; 3) sample $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$. This procedure is summarized in Algorithm 2.
Moreover, during the sampling process, we can adjust the level of stochasticity by modifying $g_{\varphi}(t)$ (eq. 12). It is important to note that changes to $g_{\varphi}(t)$ also influence $\tilde{f}^{F}_{\varphi}$ (eq. 12) and $\hat{f}_{\theta,\varphi}$ (eq. 17). In the extreme case where $g_{\varphi}(t)\equiv 0$, the reverse process becomes deterministic, allowing us to utilize off-the-shelf numerical ODE solvers and to estimate densities (Chen et al., 2018a; Grathwohl et al., 2018).
3.5 Parameterization
The parameterization of the variance function $g_{\varphi}(t)$ (eq. 12) and the data point predictor $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ (eq. 17) is straightforward. In our experiments, we parameterize them directly using neural networks. However, the parameterization of the transformation $F_{\varphi}$ (eq. 9) is more complex. NFDM necessitates that $F_{\varphi}$ be differentiable and invertible with respect to $\varepsilon$, as well as provide access to the logarithm of the determinant of the transformation $\log|J_{F}^{-1}|$ (eq. 14). Therefore, we use the following parameterization for $F_{\varphi}$:
$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t),\quad\textrm{where}$ (22)
$\mu_{\varphi}(\mathbf{x},t)$ $=(1-t)\mathbf{x}+t(1-t)\bar{\mu}_{\varphi}(\mathbf{x},t),$ (23)
$\sigma_{\varphi}(\mathbf{x},t)$ $=\delta+t\bar{\sigma}_{\varphi}\big{(}(1-t)\mathbf{x},t\big{)}/\bar{\sigma}_{\varphi}(\mathbf{0},1).$ (24)
In this context, $\bar{\mu}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ and $\bar{\sigma}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}_{+}$, with $\delta$ being a constant. In our experiments, we set $\delta^{2}=10^{-4}$. If $\bar{\mu}_{\varphi}$ and $\bar{\sigma}_{\varphi}$ are differentiable, this parameterization meets the requirements on $F_{\varphi}$ and renders $q(\mathbf{z}_{t}|\mathbf{x})$ a conditional Gaussian distribution with parameterized mean and diagonal covariance matrix.
Furthermore, this reparameterization ensures that $q(\mathbf{z}_{0}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{0};\mathbf{x},\delta^{2}I)$ and $q(\mathbf{z}_{1}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{1};0,I)$. These restrictions are not necessary for $F_{\varphi}$, however, it eliminates the need to optimize the reconstruction loss ${\mathcal{L}_{\mathrm{rec}}}$ (eq. 19) and the prior loss ${\mathcal{L}_{\mathrm{prior}}}$ (eq. 21), as they are independent of the parameters $\varphi$ and $\theta$.
The chosen parameterization of $F_{\varphi}$, $g_{\varphi}(t)$, and $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ is not the only possible approach and may not be optimal. For instance, it is possible to parameterize $F_{\varphi}$ as a flow-based model (Dinh et al., 2016; Kingma & Dhariwal, 2018), enabling a broader range of distributions than just conditional Gaussians. However, for the purpose of this paper, we maintain a simple setup and leave the exploration of the most effective parameterization for future research.
We discuss the parameterization in more detail in Section C.2.
Table 1: Comparison of NFDM results with baselines on density estimation tasks. We present results in terms of BPD, lower is better. NFDM achieves state-of-the-art results accross the three benchmark tasks.
Model CIFAR10 ImageNet 32 ImageNet 64
DDPM (Ho et al., 2020) $3.69$
Score SDE (Song et al., 2020b) $2.99$
Improved DDPM (Nichol & Dhariwal, 2021) $2.94$ $3.54$
VDM (Kingma et al., 2021) $2.65$ $3.72$ $3.40$
Score Flow (Song et al., 2021) $2.83$ $3.76$
Flow Matching (Lipman et al., 2022) $2.99$ $3.53$ $3.31$
Stochastic Interp. (Albergo & Vanden-Eijnden, 2022) $2.99$ $3.48$
i-DODE (Zheng et al., 2023b) $2.56$ $3.43$
MuLAN (Sahoo et al., 2023) $2.55$ $3.67$
NFDM (this paper) $\mathbf{2.49}$ $\mathbf{3.36}$ $\mathbf{3.20}$
4 Restricted NFDM
We introduce NFDM as a powerful framework that enables both pre-definition and learning of the forward process, potentially simplifying the task for the reverse process. However, there are in general an infinite number of forward and reverse processes that correspond to each other. So how to choose between them? The standard NFDM, as described in Section 3, learns one such pair of processes. However, the flexibility of NFDM opens up the possibility of choosing specific pairs of processes depending on the task.
Suppose our objective is not just to learn any generative process for the data, but to learn one endowed with specific beneficial properties. For instance, one such property is straight generative ODE trajectories. Learning a generative process with straight ODE trajectories can be highly beneficial, as it enables the generation of samples with far fewer steps in the ODE solver. This can lead to significant computational savings.
To achieve this, we must introduce some restrictions on the reverse process. One approach is to directly encode this in the definition of the reverse process. Alternatively, we can introduce penalties on the curvature of its trajectories. In both cases learnable forward process would then adapt to align with the reverse process. As a result, we would obtain a generative process that not only corresponds to the forward process, ensuring accurate data generation, but also features the desired property of straight trajectories. We discuss NFDM with restrictions in more details in Section A.2.
Importantly, conventional diffusion models are incapable of handling such constraints or penalization strategies. In diffusion models with a fixed forward process, the target for the reverse process is predetermined and the corresponding ODE trajectories are highly curved. Hence, imposing constraints on the reverse process, such as trajectory straightness, would lead to a mismatch with the forward process and, consequently, an inability to generate samples with high data fidelity.
4.1 NFDM with optimal transport
In this section, we discuss curvature penalization on the ODE trajectories of the reverse process in NFDM. We refer to this variant as NFDM-OT [^1] OT stands for optimal transport. This designation is used for convenience. While straight trajectories are a necessary condition for dynamic optimal transport, they are not sufficient..
We propose learning a model with an additional curvxature penalty as suggested by Kelly et al. (2020):
${\mathcal{L}_{\mathrm{OT}}}$ $=\mathcal{L}+\lambda{\mathcal{L}_{\mathrm{crv}}},\quad\textrm{where}$ (25)
${\mathcal{L}_{\mathrm{crv}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}$ $\left\|\frac{d\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{dt}\right\|_{2}^{2}.$ (26)
${\mathcal{L}_{\mathrm{crv}}}$ is an additional curvature loss that penalizes the second time derivative of the generative ODE trajectories. $\mathcal{L}$ is estimated as in section 3.3, whereas when calculating ${\mathcal{L}_{\mathrm{crv}}}$ we set $g_{\varphi}(t)\equiv 0$. ${\mathcal{L}_{\mathrm{crv}}}\equiv 0$ ensures that the generative trajectories are straight lines. Although we parameterize $\hat{f}_{\theta,\varphi}$ (eq. 17) through $\tilde{f}^{B}_{\varphi}$ (eq. 15), ${\mathcal{L}_{\mathrm{crv}}}$ does not penalize the curvature of the conditional forward trajectories. The penalty requires $F_{\varphi}$ to be twice differentiable with respect to time. The specifics of the curvature loss are elaborated upon in more detail in Section A.2.
Even though the parameterization, as discussed in section 3.5, is simple and does not guarantee reaching the zero curvature, we find that in practice it is flexible enough to successfully learn straight line generative trajectories. In our experiments, we set $\lambda=10^{-2}$. Empirical evidence suggests that higher values of $\lambda$ lead to slower convergence of the model. We provide empirical results for NFDM-OT in Section 5.2.
Table 2: Summary of FID results for few-step generation. The table is divided into three sections, based on different types of methods: those that do not minimize curvature, solvers for pretrained models, and models that specifically aim to minimize curvature. For the DDPM, we include results corresponding to two distinct objectives: the full ELBO-based objective and a simplified objective (${\mathcal{L}_{\mathrm{simple}}}$). NFDM-OT outperforms baselines with comparable NFE values.
CIFAR-10 ImageNet 32 ImageNet 64
Model NFE $\downarrow$ FID $\downarrow$ NFE $\downarrow$ FID $\downarrow$ NFE $\downarrow$ FID $\downarrow$
DDPM (${\mathcal{L}_{\mathrm{simple}}}$) (Ho et al., 2020) $1000$ $3.17$
DDPM (ELBO) (Ho et al., 2020) $1000$ $13.51$
Flow Matching (Lipman et al., 2022) $142$ $6.35$ $122$ $5.02$ $138$ $14.14$
DDIM (Song et al., 2020a) $10$ $13.36$
$12$ $5.28$
DPM Solver (Lu et al., 2022) $24$ $2.75$
Trajectory Curvature Minimization (Lee et al., 2023) $5$ $18.74$
$4$ $17.28$ $4$ $38.45$
Multisample Flow Matching (Pooladian et al., 2023) $12$ $7.18$ $12$ $17.6$
$2$ $12.44$ $2$ $9.83$ $2$ $27.70$
$4$ $7.76$ $4$ $6.13$ $4$ $17.28$
NFDM-OT (this paper) $12$ $5.20$ $12$ $4.11$ $12$ $11.58$
5 Experiments
We first showcase results demonstrating that NFDM consistently achieves better likelihood compared to baselines, obtaining state-of-the-art diffusion modeling results on the CIFAR-10 and downsampled ImageNet datasets. Then, we explore the NFDM-OT modification, which penalizes the curvature of the deterministic generative trajectories. The NFDM-OT reduces trajectory curvature, significantly lowering the number of generative steps required for sampling.
We evaluate NFDM on synthetic data, CIFAR-10 (Krizhevsky et al., 2009), and two downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) datasets. For parameterization we use multi-layer perceptrons for synthetic data and the U-Net architecture from Dhariwal & Nichol (2021) for images.
We report the NLL in Bits Per Dimension (BPD) and sample quality measured by the FID score (Heusel et al., 2017). The NLL is calculated by integrating the ODEs using the RK45 solver Dormand & Prince (1980), with all NLL metrics computed on test data. For the FID, we provide the average over 50k images.
For a detailed description of parameterizations and other experimental details, please refer to Appendix C.
(a) Score SDE (Song et al., 2020b)
(b) NFDM (this paper)
Figure 1: Comparison of trajectories between the data distribution (on the left) and the prior distribution (on the right).
5.1 Likelihood estimation
We have trained NFDM on the CIFAR-10 and downsampled ImageNet datasets. The NLL results are summarized in Table 1. Notably, NFDM outperforms diffusion-based baselines on all three datasets. We consider this enhanced performance of the NFDM a natural progression due to the following reasons.
First, it is well-established (Song et al., 2021; Zheng et al., 2023b) that diffusion models exhibit improved likelihood estimation when trained with the full ELBO objective. As delineated in Section 3.3, the objective $\mathcal{L}$ (eq. 18) used for training the NFDM constitutes a variational bound on the likelihood.
Second, as previously mentioned, diffusion models can be seen as hierarchical VAEs. From this perspective, the baselines in Table 1 resemble VAEs with either fixed or constrained variational distributions. In contrast, the NFDM extends beyond these baselines by providing a more flexible variational distribution. This flexibility allows the NFDM to better conform to the reverse process, consequently enhancing likelihood estimation.
5.2 Straight trajectories
We next evaluate NFDM-OT, which is designed to penalize the curvature of deterministic generative trajectories. Initially, we compare NFDM-OT with a conventional continuous-time diffusion model. Figure 1(a) illustrates deterministic trajectories between a two-dimensional data distribution and a unit Gaussian distribution learnt by a conventional diffusion model (Song et al., 2020b). Figure 1(b) depicts trajectories learnt by NFDM-OT. Conventional diffusion, being constrained in its forward process, learns highly curved trajectories, whereas NFDM-OT successfully learns straight generative trajectories as desired.
Subsequently, we present evaluation results of NFDM-OT on image datasets. Table 2 reports the FID scores for $2$, $4$, and $12$ Number of Function Evaluations (NFE) with respect to the function $\hat{f}_{\theta,\varphi}$ (eq. 16). In this experiment, we employ Euler’s method for sampling integration. For the specified NFEs, NFDM-OT demonstrates superior sample quality compared to other approaches with similar NFE values. Specifically, NFDM-OT outperforms approaches that minimize the curvature of generative trajectories (Pooladian et al., 2023; Lee et al., 2023).
Importantly, NFDM-OT is trained with an ELBO-based objective (eq. 18), which is known to yield higher FID scores for diffusion models (Song et al., 2021; Zheng et al., 2023b). In contrast, some of the approaches listed in Table 2 are trained with different objectives, leading to improved FID scores. Even so, for comparable NFE values NFDM-OT still achieves superior results.
The primary aim of the experiments with NFDM-OT is not to introduce a novel model that surpasses others in few-step generation, but rather to showcase NFDM’s ability to learn generative dynamics with specific properties. The straightness of the trajectories is just one example of such properties. We leave it for future research to explore different parameterizations and a modified objectives for NFDM that may yield even better results.
6 Related work
Diffusion models, originally proposed by Sohl-Dickstein et al. (2015), have evolved significantly through subsequent developments (Song & Ermon, 2019; Ho et al., 2020). These advancements have resulted in remarkable generative quality for high-dimensional data distributions (Dhariwal & Nichol, 2021; Saharia et al., 2022). Nevertheless, conventional diffusion models, typically relying on a linear Gaussian forward process, may not optimally fit some data distributions. To address this, alternative forward processes have been explored, such as combining blurring with Gaussian noise injection (Rissanen et al., 2022; Daras et al., 2022; Hoogeboom & Salimans, 2022), diffusion in the wavelet spectrum (Phung et al., 2023), and forward processes based on the exponential family (Okhotin et al., 2023). These models are limited by their fixed forward processes and may be seen as specific instances of NFDM.
Some studies have focused on making the forward process learnable. Approaches include a learnable noise injection schedule (Kingma et al., 2021; Nichol & Dhariwal, 2021; Sahoo et al., 2023) and learning data transformations like time-independent transformations based on VAEs (Vahdat et al., 2021; Rombach et al., 2022) and normalizing flows (Kim et al., 2022), or time-dependent transformations (Gu et al., 2022; Nielsen et al., 2023; Bartosh et al., 2023). These methods can also be considered special cases of NFDM with specific transformations $F_{\varphi}$ (eq. 9).
Recent studies (Zhang & Chen, 2021; De Bortoli et al., 2021; Shi et al., 2023) have explored generative models based on Schrödinger Bridge theory and finite-time diffusion constructions. While these models offer novel approaches to learning forward transformations, in contrast to NFDM, they depart from the simulation-free paradigm by requiring the full simulation of the stochastic processes for training, increasing their computational cost.
Despite the simulation-free nature of the training procedure, diffusion models still necessitate full reverse process simulations for sample generation, leading to slow and computationally expensive inference. To address this, in an orthogonal line of works alternative sampling methods have been studied, such as deterministic sampling in discrete (Song et al., 2020a) and continuous time (Song et al., 2020b) and novel numerical solvers (Tachibana et al., 2021; Liu et al., 2022a; Lu et al., 2022; Shaul et al., 2023b). Additionally, distillation techniques have been applied to both discrete (Salimans & Ho, 2022) and continuous time models (Song et al., 2023; Zheng et al., 2023a; Yin et al., 2023) to enhance sampling speed, albeit at the cost of training additional models. Most of these approches are compatible with NFDM and may be combined for further gains.
The significance of deterministic trajectories in efficient sampling was highlighted by Karras et al. (2022). At the same time, building on diffusion model concepts, Lipman et al. (2022); Liu et al. (2022b); Albergo & Vanden-Eijnden (2022) introduced simulation-free methods for learning deterministic generative dynamics. Based on these ideas, Liu et al. (2022b) and Liu (2022) proposed distillation procedures to straighten deterministic generative trajectories. Since these methods are based on distillation, we consider them orthogonal to NFDM.
Pooladian et al. (2023) and Tong et al. (2023) proposed to construct the forward process with optimal data-noise couplings to learn straighter generative trajectories. While this method is theoretically justified, it relies on estimating an optimal coupling over minibatches of the entire dataset, which, for large datasets, may become uninformative as to the true coupling. In contrast, NFDM-OT directly penalizes curvature of generative trajectories.
Lee et al. (2023) suggested learning the forward process distribution $q(\mathbf{z}_{1}|\mathbf{x})$ and interpolating linearly between $\mathbf{x}$ and $\mathbf{z}_{1}$. Similarly, Shaul et al. (2023a) aimed to modify the forward process to reduce curvature in the generative process. Both approaches can be viewed as specific instances of NFDM.
Lastly, Albergo et al. (2023) suggested optimizing the forward process simultaneously to training the reverse process, resultin in a challenging high dimentional min-max problem. Moreover, only the theoretical algorithm was presented, without empirical study and descriptions of ways to parameterize such a learnable forward process.
We continue discussing related works in Appendix B.
7 Discussion and Limitations
In this paper we introduced NFDM, a novel simulation-free framework for improved diffusion modeling through a learnable forward processes. Our approach outperforms baseline diffusion models on standard benchmarks, showcasing the effectiveness of NFDM in log-likelihood estimation.
We introduce NFDM not merely as a specific model, but rather as a versatile framework that facilitates the predefined specification and learning of the forward process. For the sake of simplicity in this work, we have chosen a straightforward Gaussian parameterization for the forward process (see Section 3.5). However, NFDM is capable of supporting far more flexible parameterizations beyond standard Gaussian distributions.
Similarly, the penalty on the curvature of deterministic trajectories discussed in Section 4.1 represents just one specific example of the restrictions that NFDM can accommodate. NFDM opens up significant prospects for exploring new generative dynamics.
Nonetheless, the advantages of NFDM come with certain trade-offs. Once the forward process is parameterized using a neural network, this leads to increased computational costs compared to conventional diffusion models. In our experiments, an optimization iteration of NFDM takes approximately 2.2 times longer than that of conventional diffusion models.
Despite these challenges, we are optimistic about the potential of NFDM in various domains and practical applications, given its flexibility in learning generative processes. We also believe that with alternative parameterizations, modifications of the objective, and the integration of orthogonal approaches like distillation, NFDM has the potential to achieve even better results.
Appendix A Derivations
A.1 Derivation of objective
For the forward process, we have the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which is implicitly defined by the invertible transformation $\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x})$ (eq. 9), where $\varepsilon\sim q(\varepsilon)$. Additionally, there’s a SDE (eq. 15) starting from $\mathbf{z}_{1}\sim q(\mathbf{z}_{1}|\mathbf{x})$ and flowing backwards in time:
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w}.$ (27)
This SDE corresponds to the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
The reverse process involves a prior distribution $p(\mathbf{z}_{1})$ and a reverse SDE (eq. 15):
$d\mathbf{z}_{t}=$ $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}}.$ (28)
Our goal is to derive a variational bound on the negative log-likelihood of the model.
We begin by discretizing the conditional reverse SDE in eq. 27 and the reverse SDE in eq. 28, transitioning from continuous-time trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ to discrete-time trajectories $\bar{\mathbf{z}}_{0},\bar{\mathbf{z}}_{\frac{1}{T}},\dots,\bar{\mathbf{z}}_{1}$, where $T$ represents the number of discrete steps. For convenience, we also introduce $\Delta t=\frac{1}{T}$.
The discretization of the conditional reverse SDE is as follows:
$\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})=q_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x}),\quad\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x}),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (29)
We similarly discretize the reverse SDE:
$\bar{p}(\bar{\mathbf{z}}_{1})=p(\bar{\mathbf{z}}_{1}),\quad\bar{p}(\mathbf{x}|\bar{\mathbf{z}}_{0})=p(\mathbf{x}|\bar{\mathbf{z}}_{0}),\quad\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (30)
It’s important to note that once we discretize the trajectories, $\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t}|\mathbf{x})\neq q_{\varphi}(\bar{\mathbf{z}}_{t}|\mathbf{x})$ and $\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t})\neq p_{\theta,\varphi}(\bar{\mathbf{z}}_{t})$ except for $t=1$. However, this discretization corresponds to an Euler-Maruyama method of integrating SDEs. Thus, we know that as $\Delta t\rightarrow 0$, the discretized trajectories converge to the continuous ones.
Now, considering these discrete processes and recognizing that they are Markovian conditional and unconditional processes moving backwards in time, we can leverage a key result from Ho et al. (2020):
$\mathbb{E}_{q(\mathbf{x})}\left[-\log\bar{p}_{\theta,\varphi}(\mathbf{x})\right]\leq\bar{\mathcal{L}}_{\mathrm{recon}}+\bar{\mathcal{L}}_{\mathrm{diff}}+\bar{\mathcal{L}}_{\mathrm{prior}},\quad\textrm{where}$ (31)
¯$\bar{\mathcal{L}}_{\mathrm{recon}}=\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{0})}\left[-\log p(\mathbf{x}|\bar{\mathbf{z}}_{0})\right],$ (32)
¯$\bar{\mathcal{L}}_{\mathrm{diff}}=\sum_{t=\frac{1}{T}}^{T}\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{t})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})\|\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})\big{)}\right],$ (33)
¯$\bar{\mathcal{L}}_{\mathrm{prior}}=\mathbb{E}_{q(\mathbf{x})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})\|\bar{p}(\bar{\mathbf{z}}_{1})\big{)}\right].$ (34)
Consequently, eq. 31 establishes a variational upper bound on the negative log-likelihood of discretized reverse process. Taking the limits of this upper bound, the reconstruction term ¯$\bar{\mathcal{L}}\mathrm{recon}$ and the prior term ¯$\bar{\mathcal{L}}\mathrm{prior}$ retain their forms. However, we can reformulate the diffusion term:
$\mathcal{L}_{\mathrm{diff}}$ $=\lim_{\Delta t\rightarrow 0}\bar{\mathcal{L}}_{\mathrm{diff}}$ (35)
$=\lim_{\Delta t\rightarrow 0}\sum_{t=\frac{1}{T}}^{T}\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{t})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})\|\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})\big{)}\right]$ (36)
[…]
$=\mathbb{E}_{u(t)\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{t})}\left[\frac{1}{2g^{2}_{\varphi}(t)}\left\|\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t)\right\|_{2}^{2}\right],$ (39)
where $u(t)$ is a uniform distribution over a unit interval.
Therefore, as the discretized trajectory tends to the continuous one when $\Delta t\rightarrow 0$, we can use the obtained limits as a variational bound on the negative log-likelihood of the continuous model.
A.2 NFDM with restrictions
In Section 4, we discussed training the NFDM with restrictions on the reverse process. This section aims to explore the limitations of such restrictions and to further discuss the training of NFDM with curvature penalties.
When training NFDM with certain restrictions, it is crucial that these restrictions are feasible. For example, imposing overly stringent constraints on the reverse process (such as fixing its drift term $\hat{f}\equiv 0$) will render the forward process incapable of adapting, regardless of its flexibility. Similarly, unattainable penalties, like those on the squared norm of the drift term $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}$, will lead to biased solutions, as it’s impossible to match two processes when $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}\equiv 0$.
Hence, in Section 4.1, we suggest penalizing the curvature of deterministic generative trajectories specifically with penalty ${\mathcal{L}_{\mathrm{crv}}}$ (eq. 26). This is because for non-degenerate data distributions, there exist mappings that transform them into unit Gaussian noise along straight line trajectories.
To calculate the curvature penalty ${\mathcal{L}_{\mathrm{crv}}}$ (eq. 26), we proceed as follows. First, we use the chain rule to rewrite the time derivative:
$\frac{d\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{dt}$ $=\frac{\partial\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{\partial\mathbf{z}_{t}}\frac{\partial\mathbf{z}_{t}}{\partial t}+\frac{\partial\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{\partial t}$ (40)
$=\frac{\partial\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{\partial\mathbf{z}_{t}}\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)+\frac{\partial\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{\partial t}.$ (41)
The second term in eq. 41 is the time derivative of a function. As discussed in section 3.1, the time derivatives can be determined as a Jacobian Vector Product (JVP)with a one-dimensional unit vector. Notably, the first term in eq. 41 is also a JVP. Therefore, we can combine these two operations. For this purpose, we define a $\mathbb{R}^{D+1}$ dimensional adjoint vector $v$:
$v=\begin{bmatrix}\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\\ 1\end{bmatrix}.$ (42)
Consequently, ${\mathcal{L}_{\mathrm{crv}}}$ can be computed as the JVP of $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)$ with the vector $v$.
Appendix B Connections of NFDM with other approaches
B.1 Special cases of NFDM
Many existing works define the forward process in diffusion models as either fixed or simply parameterized processes. These processes can be considered special cases of NFDM. In this section, we review some of these approaches.
Soft Diffusion. Daras et al. (2022) consider a more general case of fixed forward processes, $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};C_{t}\mathbf{x},\sigma_{t}I)$, which can be parameterized as $F(\varepsilon,t,\mathbf{x})=C_{t}\mathbf{x}+\sigma_{t}\varepsilon$. Such distributions include, for example, combinations of blurring and the injection of Gaussian noise.
Star-Shaped Diffusion. Okhotin et al. (2023) extended conventional diffusion models to include distributions from the exponential family. Although Star-Shaped Diffusion is a discrete-time approach and does not directly correspond to NFDM, the latter can work with exponential family distributions through reparameterization functions. For instance, for some continuous one-dimensional distribution $q(\mathbf{z}_{t}|\mathbf{x})$, NFDM could use $F(\varepsilon,t,\mathbf{x})=a(b(\varepsilon),t,\mathbf{x})$, where $a$ is the inverse Cumulative Distribution Function (CDF) of $q(\mathbf{z}_{t}|\mathbf{x})$, and $b$ is the CDF of a unit Gaussian.
Variational Diffusion Models. Kingma et al. (2021) proposed forward conditional distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ as $\mathcal{N}(\mathbf{z}_{t};\alpha_{\varphi}(t)\mathbf{x},\sigma_{\varphi}^{2}(t)I)$ with learnable parameters $\varphi$. In the context of NFDM, this distribution can be parameterized by $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{\varphi}(t)\mathbf{x}+\sigma_{\varphi}(t)\varepsilon$.
LSGM. Vahdat et al. (2021) suggest an alternative approach for parameterizing the forward process, proposing diffusion in a latent space of a VAE. Therefore, the forward process is characterized by a distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}E_{\varphi}(\mathbf{x}),\sigma_{t}^{2}I)$, where $E_{\varphi}$ is the encoder of the VAE. To parameterize the same forward process with NFDM, one could use $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}E_{\varphi}(\mathbf{x})+\sigma_{t}\varepsilon$. To align the reverse process, the reconstruction distribution should be $p(\mathbf{x}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{x};D_{\varphi}(\mathbf{z}_{0}),\delta^{2}I)$, where $D_{\varphi}$ is VAE’s decoder.
DiffEnc. Nielsen et al. (2023) proposed a more general forward process, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}f_{\varphi}(\mathbf{x},t),\sigma_{t}^{2}I)$, which, unlike LSGM, transforms $\mathbf{x}$ in a time-dependent manner. This forward process can also be described in terms of NFDM as $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}f_{\varphi}(\mathbf{x},t)+\sigma_{t}\varepsilon$.
B.2 Connections of NFDM objective
In this section, we delve into the details of the NFDM’s objective function (eq. 18), with a particular focus on the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ (eq. 20).
First, let’s unpack the ${\mathcal{L}_{\mathrm{diff}}}$ by substituting the definitions of $\tilde{f}^{B}_{\varphi}$ and $\hat{f}_{\theta,\varphi}$:
${\mathcal{L}_{\mathrm{diff}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\right\|_{2}^{2}\right]$ (43)
$=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\underbrace{\Big{(}f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\Big{)}}_{\textrm{Flow Matching term}}+\frac{g_{\varphi}^{2}(t)}{2}\underbrace{\Big{(}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\Big{)}}_{\textrm{Score Matching term}}\right\|_{2}^{2}\right].$ (44)
This formulation clearly delineates two components of the objective: the first calculates the difference between the ODE drift terms, and the second calculates the difference between the score functions. Moreover, this expression highlights the role of $g_{\varphi}(t)$. When $g_{\varphi}(t)$ is small, the forward and reverse processes exhibit smoother trajectories, and the objective is dominated by the first term. Conversely, when $g_{\varphi}(t)$ is large, the processes exhibit more stochastic trajectories, and the objective is dominated by the second term.
Crucially, in the extreme scenarios where $g_{\varphi}(t)$ approaches either 00 or $\infty$, the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ corresponds to either a reweighted Flow Matching loss (Lipman et al., 2022) or a reweighted Score Matching loss (Vincent, 2011), respectively:
$\lim_{g_{\varphi}(t)\rightarrow 0}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow 0}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\right\|_{2}^{2}\right],$ (45)
$\lim_{g_{\varphi}(t)\rightarrow\infty}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow\infty}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{g_{\varphi}^{2}(t)}{4}\left\|\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\right\|_{2}^{2}\right]$ (46)
B.3 Connections of NFDM with Stochastic Interpolants
In this section, we explore the connections between NFDM and Stochastic Interpolants, as proposed by Albergo et al. (2023). Both NFDM and Stochastic Interpolants introduce a more general family of forward processes through the use of a reparameterization function. Although we acknowledge the relevance of Stochastic Interpolants to our work, there are notable differences between the two approaches.
First, our methodology involves parameterizing the reverse process by substituting the prediction of $\mathbf{x}$ into the forward process, whereas Stochastic Interpolants necessitate learning two separate functions for the reverse process: the velocity field and the score function.
Second, we present NFDM as a framework that enables both the pre-specification and learning of the forward process, in contrast to Stochastic Interpolants, which are derived under the assumption of a fixed forward process. Consequently, the objectives utilized by Stochastic Interpolants do not support the incorporation of learnable parameters for the forward process. Furthermore, these objectives are tailored towards learning the velocity field and the score function, rather than optimizing likelihood, as is the case with NFDM.
In their practical applications, Stochastic Interpolants are demonstrated with only simple parameterizations of the forward process. They propose a theoretical framework for learning the forward process that would result in an reverse process characterized by dynamical optimal transport. However, this approach is contingent upon solving a high-dimensional min-max optimization problem, for which they do not provide experimental results. Moreover, their work does not clearly articulate how Stochastic Interpolants might be applied to learning other generative dynamics.
In contrast, NFDM introduces a more generalized method for learning generative dynamics. Moreover, when NFDM is learned with restrictions or penalties (see Section 4), it remains within a min-min optimization paradigm, which can be addressed more efficiently.
Appendix C Implementation details
Our evaluation of NFDM includes tests on synthetic data, CIFAR-10 (Krizhevsky et al., 2009), and two downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) datasets. To maintain consistency with baselines, we employ horizontal flipping as a data augmentation technique in training models on CIFAR-10 and ImageNet (Song et al., 2020b, 2021). For density estimation of discrete data, uniform dequantization is used (see Section C.1).
We parameterize $\hat{\mathbf{x}}_{\theta}$ (eq. 17) in the reverse process using a 5-layer perceptrons with 512 neurons in each layer for synthetic data and the U-Net architecture from Dhariwal & Nichol (2021) for images. In all experiments, a 3-layer perceptrons with 64 neurons in each layer is employed to parameterize $g_{\varphi}$ (eq. 12), and for $F_{\varphi}$ (eq. 9), we use an identical neural network to that of $\hat{\mathbf{x}}_{\theta}$. The sole difference is that for $F_{\varphi}$, we double the output of the last layer to parameterize $\bar{\mu}_{\varphi}$ (eq. 23) and $\bar{\sigma}_{\varphi}$ (eq. 24) with same model (see Section 3.5).
The models were trained using the Adam optimizer with the following parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, a weight decay of $0.0$, and $\epsilon=10^{-8}$. The training process was facilitated by a polynomial decay learning rate schedule, which includes a warm-up phase for a predefined number of training steps. During this phase, the learning rate is linearly increased from $10^{-8}$ to a peak value. After reaching the peak learning rate, it is then linearly decreased to $10^{-8}$ by the final training step. The specific hyperparameters are detailed in Table 3. Training was carried out on Tesla V100 GPUs
Table 3: Training hyper-parameters.
CIFAR-10 ImageNet 32 ImageNet 64
Channels 256 256 192
Depth 2 3 3
Channels multipliers 1,2,2,2 1,2,2,2 1,2,3,4
Heads 4 4 4
Heads Channels 64 64 64
Attention resolution 16 16,8 32,16,8
Dropout 0.0 0.0 0.0
Effective Batch size 256 1024 2048
GPUs 2 4 16
Epochs 1000 200 250
Iterations 391k 250k 157k
Learning Rate 4e-4 1e-4 1e-4
Learning Rate Scheduler Polynomial Polynomial Constant
Warmup Steps 45k 20k -
C.1 Dequantization
For reporting the NLL, we employ standard uniform dequantization. The NLL is estimated using an importance-weighted average, given by
$\log\frac{1}{K}\sum_{k=1}^{K}p_{\theta}(\mathbf{x}+u_{k}),\quad\text{where}\quad u_{k}\sim\mathcal{U}(0,1),$ (47)
where $\mathbf{x}\in[0,\dots,255]$.
C.2 Parameterization
We define the parameterization of the function $F_{\varphi}$ as follows in eq. 22:
$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t).$ (48)
In addition to inferring $F_{\varphi}$, we require access to its inverse $F^{-1}_{\varphi}$ and the logarithm of the determinant of the inverse transformation $\log|J_{F}^{-1}|$ (eq. 14). The calculation of the inverse is straightforward:
$F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=\frac{\mathbf{z}_{t}-\mu_{\varphi}(\mathbf{x},t)}{\sigma_{\varphi}(\mathbf{x},t)}.$ (49)
The logarithm of the determinant of the inverse transformation, $\log|J_{F}^{-1}|$, can also be readily determined:
$\log|J_{F}^{-1}|=-\log|J_{F}|=-\sum_{i=1}^{D}\log\left[\sigma_{\varphi}(\mathbf{x},t)\right]_{i},$ (50)
where $D$ represents the dimensionality of the data, and $\left[\sigma_{\varphi}(\mathbf{x},t)\right]_{i}$ refers to $i$-th output of the $\sigma_{\varphi}$ function.
To ensure the constraints $g_{\varphi}\geq 0$ (eq. 12) and $\bar{\sigma}_{\varphi}\geq 0$ (eq. 24) are met, we apply the softplus function to the outputs of the neural networks.

### v2

@@
 Abstract
-Conventional diffusion models typically relies on a fixed forward process, which implicitly defines complex marginal distributions over latent variables.
+Conventional diffusion models often rely on a fixed forward process, which implicitly defines complex marginal distributions over latent variables.
 This can often complicate the reverse process’ task in learning generative trajectories, and results in costly inference for diffusion models.
-To address these limitations, we introduce Neural Flow Diffusion Models (NFDM), a novel framework that enhances diffusion models by supporting a broader range of forward processes beyond the fixed linear Gaussian.
+To address these limitations, we introduce Neural Flow Diffusion Models (NFDM), a novel framework that enhances diffusion models by supporting a broader range of forward processes beyond the standard linear Gaussian.
 We also propose a novel parameterization technique for learning the forward process.
 Our framework provides an end-to-end, simulation-free optimization objective, effectively minimizing a variational upper bound on the negative log-likelihood.
-Experimental results demonstrate NFDM’s strong performance, evidenced by state-of-the-art likelihood estimation.
-Furthermore, we investigate NFDM’s capacity for learning generative dynamics with specific characteristics, such as deterministic straight lines trajectories.
-This exploration underscores NFDM’s versatility and its potential for a wide range of applications.
-diffusion, generative models, variational inference
+Experimental results demonstrate NFDM’s strong performance, evidenced by state-of-the-art likelihoods across a range of image generation tasks.
+Furthermore, we investigate NFDM’s capacity for learning generative dynamics with specific characteristics, such as deterministic straight lines trajectories, and demonstrate how the framework can be adopted for learning bridges between two distributions.
+The results underscores NFDM’s versatility and its potential for a wide range of applications.
 1 Introduction
-Generative models are a versatile class of probabilistic machine learning models, which finds applications across diverse fields such as art, music, medicine, and physics (Tomczak, 2022; Creswell et al., 2018; Papamakarios et al., 2021; Yang et al., 2022).
-Generative models excel in constructing probability distributions that accurately represent datasets, enabling them to produce new samples akin to the original data.
-Such capabilities make them ideal for tasks like enhancing training data through the generation of synthetic datasets (data augmentation) and uncovering hidden patterns and structures in an unsupervised learning setting.
-Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of generative models are constructed by two key processes: the forward process and the reverse process.
+Diffusion models [55, 19] are a class of generative models constructed by two key processes: the forward process and the reverse process.
 The forward process gradually corrupts the data distribution, transforming it from its original form to a noised state.
 The reverse process learns to invert corruptions of the forward process and restore the data distribution.
-This way, it learns to generate data from pure noise.
-Diffusion models have demonstrated remarkable results in various domains (Ho et al., 2022; Saharia et al., 2022; Tan et al., 2024; Watson et al., 2023; Trippe et al., 2023).
-Nevertheless, most existing diffusion models fix the forward process to be predefined Gaussian, which makes it unable to adapt to the task at hand or simplify the target for the reverse process.
-At the same time there are many works that demonstrate how modifications of the forward process improve performance in terms of generation quality (Nichol & Dhariwal, 2021; Vahdat et al., 2021; Daras et al., 2022), likelihood estimation (Kingma et al., 2021; Nielsen et al., 2023; Bartosh et al., 2023) or sampling speed (Lee et al., 2023; Pooladian et al., 2023; Tong et al., 2023).
-In this paper, we present Neural Flow Diffusion Models (NFDM), a framework that allows for the pre-specification and learning of latent variable distributions defined by the forward process.
-Unlike conventional diffusion models (Ho et al., 2020), which rely on a conditional Gaussian forward process, NFDM may accommodate any continuous (and learnable) distribution that can be expressed as an invertible mapping applied to noise.
-We also derive, and leverage, an end-to-end simulation-free optimization procedure, that minimizes a variational upper bound on the negative log-likelihood (NLL).
-We also propose an efficient neural network-based parameterization for the forward process, enabling it to adapt to the reverse process during training and simplify the learning of the data distribution.
-To demonstrate NFDM’s capabilities with a learnable forward process we provide experimental results on CIFAR-10, ImageNet 32 and 64, attaining state-of-the-art NLL results, which is crucial for many applications such as data compression (Ho et al., 2021; Yang & Mandt, 2022), anomaly detection (Chen et al., 2018b; Dias et al., 2020) and out-of-distribution detection (Serrà et al., 2019; Xiao et al., 2020).
-Leveraging the flexibility of NFDM, we further explore training with constraints on the reverse process to learn generative dynamics with specific properties.
-As a case study, we discuss a curvature penalty on the deterministic generative trajectories.
-Our empirical results indicate improved computational efficiency compared to baselines on synthetic datasets as well as MNIST, CIFAR-10, and downsampled ImageNet.
+This way, the model learns to generate data from pure noise.
+Diffusion models have demonstrated remarkable results in various domains [20, 47, 62, 68, 64].
+Nevertheless, most existing diffusion models fix the forward process to be predefined, usually linear, Gaussian which makes it unable to adapt to the task at hand or simplify the target for the reverse process.
+At the same time there is a growing body of work that demonstrates how modifications of the forward process improve performance in terms of generation quality [38, 65, 9], likelihood estimation [27, 39, 4] or sampling speed [31, 43, 63].
+In this paper, we present Neural Flow Diffusion Models (NFDM), a framework that allows for the pre-specification and learning of complex latent variable distributions defined by the forward process.
+Unlike conventional diffusion models [19], which rely on a conditional Gaussian forward process, NFDM can accommodate any continuous (and learnable) distribution that can be expressed as an invertible mapping applied to noise.
+We also derive, and leverage, a new end-to-end simulation-free optimization procedure for NFDM, that minimizes a variational upper bound on the negative log-likelihood (NLL).
+Furthermore, we propose an efficient neural network-based parameterization for the forward process, enabling it to adapt to the reverse process during training and simplify the learning of the data distribution.
+To demonstrate NFDM’s capabilities with a learnable forward process we provide experimental results on CIFAR-10, ImageNet 32 and 64, attaining state-of-the-art NLL results, which is crucial for many applications such as data compression [21, 70], anomaly detection [7, 13] and out-of-distribution detection [50, 69].
+Then, leveraging the flexibility of NFDM, we demonstrate how this framework can be applied to learn bridges between two distributions using the AFHQ dataset.
+Finally, we explore training with constraints on the reverse process to learn generative dynamics with specific properties.
+As a case study, we discuss curvature and obstacle avoidance penalties on the deterministic generative trajectories.
+Our empirical results indicate improved computational efficiency compared to baselines on CIFAR-10, downsampled ImageNet, and synthetic data.
 We summarize our contributions as follows:
@@
   2. 2.
-    We develope an end-to-end optimization procedure that minimizes an upper bound on the negative log-likelihood in a simulation-free manner.
+    We develop an end-to-end optimization procedure that minimizes an upper bound on the negative log-likelihood in a simulation-free manner.
   3. 3.
@@
   4. 4.
-    We show how NFDM can be used in learning generative processes with specific properties, exemplified by dynamics with straight line trajectories, where NFDM leads to significantly faster sampling speeds and enhanced generation quality with fewer sampling steps.
+    We show how NFDM can be used in learning bridges and generative processes with specific properties, such as dynamics with straight line trajectories, leading to significantly faster sampling speeds and enhanced generation quality with fewer sampling steps.
 2 Background
@@
 Typically, the conditional distributions evolve from some low variance distribution $q(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{z}_{0}-\mathbf{x})$ to a unit Gaussian $q(\mathbf{z}_{1}|\mathbf{x})\approx\mathcal{N}(\mathbf{z}_{1};0,I)$.
-This forward process is then reversed by starting from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and following the reverse SDE (Anderson, 1982):
-$d\mathbf{z}_{t}$ $=\tilde{f}^{B}(\mathbf{z}_{t},t)dt+g(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (2)
-$\tilde{f}^{B}(\mathbf{z}_{t},t)$ $=\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}).$ (3)
+This forward process is then reversed by starting from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and following the reverse SDE [3]:
+$d\mathbf{z}_{t}=\tilde{f}^{B}(\mathbf{z}_{t},t)dt+g(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\tilde{f}^{B}(\mathbf{z}_{t},t)=\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}).$ (2)
 Here, $\bar{\mathbf{w}}$ denotes a standard Wiener process where time flows backwards.
-Diffusion models approximate this reverse process by learning $\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t})$, known as the score function, through a $\lambda_{t}$-weighted score matching loss:
-$\underset{u(t)q(\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t})\big{\|}_{2}^{2}\right],$ (4)
+Diffusion models approximate this reverse process by learning $\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t})$, known as the score function, through a $\lambda_{t}$-weighted denoising score matching loss:
+$\underset{u(t)q(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}|\mathbf{x})\big{\|}_{2}^{2}\right],$ (3)
 where $u(t)$ represents a uniform distribution over the interval $[0,1]$, and $s_{\theta}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ is a learnable approximation.
-This loss can be reformulated as a more tractable denoising score matching loss Vincent (2011):
-$\underset{u(t)q(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\lambda_{t}\big{\|}s_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\big{\|}_{2}^{2}\right].$ (5)
-With a learned score function $s_{\theta}(\mathbf{z}_{t},t)$, one can generate a sample from the reverse process by first sampling from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and then simulating the reverse SDE:
-$d\mathbf{z}_{t}$ $=\big{[}\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)s_{\theta}(\mathbf{z}_{t},t)\big{]}dt+g(t)d\bar{\mathbf{w}},$ (6)
-resulting a sample $\mathbf{z}_{0}\sim p_{\theta}(\mathbf{z}_{0})\approx q(\mathbf{z}_{0})\approx q(\mathbf{x})$.
+With a learned score function $s_{\theta}(\mathbf{z}_{t},t)$, one can generate a sample from the reverse process by first sampling from the prior $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{z}_{1};0,I)$, and then simulating the reverse SDE, resulting in a sample $\mathbf{z}_{0}\sim p_{\theta}(\mathbf{z}_{0})\approx q(\mathbf{z}_{0})\approx q(\mathbf{x})$:
+$d\mathbf{z}_{t}$ $=\big{[}\tilde{f}^{F}(\mathbf{z}_{t},t)-g^{2}(t)s_{\theta}(\mathbf{z}_{t},t)\big{]}dt+g(t)d\bar{\mathbf{w}}.$ (4)
 Diffusion models possess several important properties.
-For example, for specific $\lambda_{t}$, the objective (eq.
-5) can be reformulated (Song et al., 2021) as an Evidence Lower Bound (ELBO) on the model’s likelihood.
+For example, for a specific $\lambda_{t}$, the objective (eq.
+3) can be reformulated [58] as an Evidence Lower Bound (ELBO) on the model’s likelihood.
 Furthermore, the minimization of denoising score matching (eq.
-5) is a simulation-free procedure.
+3) is a simulation-free procedure.
 This means that simulating either the forward or reverse processes through its SDE is not necessary for sampling $\mathbf{z}_{t}$, nor is it necessary for estimating the gradient of the loss function.
@@
 This implies that we can sample from diffusion models deterministically, allowing the use of off-the-shelf numerical ODE solvers for sampling, which may improve the sampling speed compared to stochastic sampling that requires simulating an SDE.
-Additionally, deterministic sampling enables us to compute densities by treating the model as a continuous normalizing flow, as detailed in Chen et al.
-(2018a); Grathwohl et al.
-(2018).
+Additionally, deterministic sampling enables us to compute densities by treating the model as a continuous normalizing flow, as detailed in [6, 16].
 3 Neural Flow Diffusion Models
-Diffusion models can be viewed as a specific type of hierarchical Variational Autoencoders (VAEs) (Kingma & Welling, 2013; Rezende et al., 2014), where the latent variables are inferred through a pre-specified scaling of the data point and injection of Gaussian noise.
-However, this formulation limits diffusion models in terms of the flexibility of their latent space, and makes learning of the reverse process more challenging.
-To address this limitation, we propose a generalized form of data and noise transformations that enables the definition and learning of a broad range of distributions in the latent space beyond the conventional Gaussian.
+In diffusion models the forward process defines stochastic conditional trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ and the reverse process tries to match the marginal distribution of trajectories.
+This construction can be viewed as a specific type of hierarchical Variational Autoencoders (VAEs) [29, 44].
+However, in conventional diffusion models the latent variables are inferred through a pre-specified linear combination of the data point and Gaussian noise.
+This formulation limits diffusion models in terms of the flexibility of their latent space, and makes learning of the reverse process more challenging.
+To address this limitation, we propose a generalized form of the forward process that enables the definition and learning of a broad range of distributions in the latent space.
+From a practical perspective, a more flexible forward process can simplify the task of learning the reverse process.
+From a theoretical perspective, learning of the forward process is analogous to learning the variational distribution in a hierarchical VAE, which gives a tighter bound on model’s NLL (see an extended discussion in Section B.1).
 In this section, we introduce Neural Flow Diffusion Models (NFDM)– a framework that generalizes conventional diffusion models.
-The key idea in NFDM is to define the forward process implicitly via a learnable transformation $F_{\varphi}(\varepsilon,t,\mathbf{x})$.
-This lets the user define a broad range of continuous time and data dependent forward process.
+The key idea in NFDM is to define the forward process’ conditional SDE implicitly via a learnable transformation $F_{\varphi}(\varepsilon,t,\mathbf{x})$ that defines the marginal distributions.
+This lets the user define a broad range of continuous time- and data-dependent forward processes, that the reverse process will learn to invert.
 Importantly, NFDM retains crucial properties of conventional diffusion models, like likelihood-based and simulation-free training.
 Previous diffusion models emerge as special cases when the data transformation is linear, time-independent, and/or additive Gaussian.
-The connections between NFDM and existing models are discussed in Section 6.
 3.1 Forward Process
-The forward process defines a target for the reverse process. It describes the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ in the latent space that the reverse process tries to match.
 We approach the forward process constructively.
-First, we define a conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for $t\in[0,1]$.
-Then, we introduce the corresponding conditional ODE that together with the initial distribution $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ matches the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
-Finally, we define a conditional SDE that implicitly defines distribution over trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, with the same marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
+The ultimate goal is a learnable distribution over trajectories, $\{\mathbf{z}(t)\}_{t\in[0,1]}$ given $\mathbf{x}$, realized by a conditional SDE constructed in three steps.
+First, we (implicitly) define the conditional marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for $t\in[0,1]$ using $F_{\varphi}(\varepsilon,t,\mathbf{x})$.
+Then, we introduce the corresponding conditional ODE that together with an initial distribution $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ matches the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
+Finally, we define a conditional SDE that defines a distribution over trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, with marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ for each $t$.
 Forward Marginal Distribution.
-We characterize the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ of the forward process trough a function of injected noise $\varepsilon$, time $t$, and the data $\mathbf{x}$:
+We characterize the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ of the forward process trough a function that transforms a noise sample $\varepsilon$ into $\mathbf{z}_{t}$, conditional on the time step $t$ and data point $\mathbf{x}$:
 $\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x}),$ (7)
 where $F_{\varphi}:\mathbb{R}^{D}\times[0,1]\times\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$ and $\varepsilon\sim q(\varepsilon)=\mathcal{N}(\varepsilon;0,I)$.
-This function transforms a sample from a distribution $q(\varepsilon)$ into $\mathbf{z}_{t}$, conditional on the data point $\mathbf{x}$ and time step $t$, implicitly defining the conditional distribution of latent variables $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
+This defines the conditional distribution of latent variables $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
 Additionally, eq.
-9 facilitates direct and efficient sampling from $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ through $F_{\varphi}$.
+7 facilitates direct and efficient sampling from $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ through $F_{\varphi}$ in eq.
+7.
 Conditional ODE.
-Assuming that $F_{\varphi}$ is differentiable with respect to $\varepsilon$ and $t$ and invertible with respect to $\varepsilon$, fixing specific values of $\mathbf{x}$ and $\varepsilon$ and varying $t$ from 00 to $1$ results in a smooth trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}_{1}$.
+We assume that $F_{\varphi}$ is differentiable with respect to $\varepsilon$ and $t$, invertible with respect to $\varepsilon$.
+Further, we assume that fixing specific values of $\mathbf{x}$ and $\varepsilon$ and varying $t$ from 00 to $1$ results in a smooth trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}_{1}$.
 Differentiating these trajectories over time yields a velocity field corresponding to the conditional distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, thereby defining a conditional ODE:
-$d\mathbf{z}_{t}$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt,\quad\textrm{where}$ (10)
-$f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=\left.\frac{\partial F_{\varphi}(\varepsilon,t,\mathbf{x})}{\partial t}\right|_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}.$ (11)
-Therefore, if we sample $\mathbf{z}_{0}\sim q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ and solve the ODE in eq.
-10 until time $t$, we have a sample $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
-The time derivative of $F_{\varphi}$ may be calculated using automatic differentiation tools like PyTorch (Paszke et al., 2017) or JAX (Bradbury et al., 2018).
-Specifically, we can compute a Jacobian Vector Product (JVP) with a unit one-dimensional vector.
-Unlike standard backpropagation, which corresponds to a Vector Jacobian Product (VJP), JVP does not necessitate $\mathcal{O}(D)$ backpropagation passes, enhancing scalability.
+$d\mathbf{z}_{t}=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt,\quad\textrm{where}\quad f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})=\left.\frac{\partial F_{\varphi}(\varepsilon,t,\mathbf{x})}{\partial t}\right|_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}.$ (8)
+Therefore, if we sample $\mathbf{z}_{0}\sim q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$ and solve eq. 8 until time $t$, we have $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
+The time derivative of $F_{\varphi}$ may be calculated efficiently using automatic differentiation tools like PyTorch [41] or JAX [5] (see details in Section B.2).
 Conditional SDE.
-The function $F_{\varphi}$ and the distribution of the noise $q(\varepsilon)$ together defines $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})$.
+The function $F_{\varphi}$ and the distribution of the noise $q(\varepsilon)$ together defines $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
 To completely define the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, we introduce a conditional SDE that starts from sample $\mathbf{z}_{0}$ and runs forward in time.
-With access to both the ODE and score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, the SDE (Song et al., 2020b) corresponding to the distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, is:
+With access to both the ODE and score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, an SDE [60] with marginal distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, is:
 $d\mathbf{z}_{t}=\tilde{f}^{F}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w},\quad\textrm{where}$ (9)
@@
 Here, $g_{\varphi}:[0,1]\mapsto\mathbb{R}_{+}$ is a scalar function, and $\mathbf{w}$ represents a standard Wiener process.
-It is important to note, that $g_{\varphi}$ only influences the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$, not the marginal distributions.
-The Conditional SDE in eq.
-12 has the same conditional marginal distribution, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, for any $g_{\varphi}$.
-The score function of $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is:
-$\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\nabla_{\mathbf{z}_{t}}\left[\log q(\varepsilon)+\log\left|J_{F}^{-1}\right|\right]$ (13)
-$\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})\quad\textrm{and}\quad J_{F}^{-1}=\frac{\partial F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})}{\partial\mathbf{z}_{t}}.$ (14)
-While the log-determinant of the Jacobian matrix $J_{F}^{-1}$ is not generally available in explicit form, it can be analytically evaluated for certain transformations or specific architectures of the function $F_{\varphi}$, like RealNVP style architectures (Dinh et al., 2016; Kingma & Dhariwal, 2018).
-The parameterization of $F_{\varphi}$ is further discussed in Section 3.5.
-Given access to the log probability $\log q(\varepsilon)$ and the log-determinant of $J_{F}^{-1}$, automatic differentiation tools enable the calculation of gradients and the score function.
-Unlike the time derivative, regular backpropagation can be utilized here, ensuring computationally efficient score function evaluation.
-3.2 Reverse process
+Note that $g_{\varphi}$ only influences the distribution of trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$.
+The marginal distributions are the same, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, for any choice of $g_{\varphi}$.
+The SDE in eq.
+9 requires access to the conditional score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which in the general case can be computationally costly.
+However, for $F_{\varphi}$ that allow efficient evaluation of the log-determinant of the Jacobian matrix, the score function can be calculated efficiently.
+Examples of such $F_{\varphi}$ are functions linear with respect to $\varepsilon$ or RealNVP style architectures [14, 28].
+The calculation of this score function is further discussed in Section B.3.
+3.2 Reverse (Generative) Process
 To define the reverse (generative) process, we specify a reverse SDE that starts from $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$ and runs backwards in time.
 To do so we first introduce a conditional reverse SDE that reverses the conditional forward SDE (eq.
-12).
-Following (Anderson, 1982), we define:
-$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w},\quad\textrm{where}$ (15)
+9).
+Following [3], we define:
+$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (10)
 $\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\frac{g_{\varphi}^{2}(t)}{2}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x}).$
-Secondly, we reference the reverse SDE (see eq. 15) and incorporating the prediction of $\mathbf{x}$:
-$d\mathbf{z}_{t}=$ $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (16)
-$\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)},$ (17)
+Secondly, leveraging this reverse SDE (see eq. 10) we incorporate a prediction of $\mathbf{x}$:
+$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)},$ (11)
 and $\hat{\mathbf{x}}_{\theta}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ is a function that predicts the data point $\mathbf{x}$.
 This SDE defines the dynamics of the generative trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$.
-However, to fully specify the reverse process, it is necessary to define a prior distribution $p(\mathbf{z}_{1})$ and a reconstruction distribution $p(\mathbf{x}|\mathbf{z}_{0})$.
+To fully specify the reverse process, it is also necessary to define a prior distribution $p(\mathbf{z}_{1})$ and a reconstruction distribution $p(\mathbf{x}|\mathbf{z}_{0})$.
 In all of our experiments, we set the prior $p(\mathbf{z}_{1})$ to be a unit Gaussian distribution $\mathcal{N}(\mathbf{z}_{1};0,I)$ and let $p(\mathbf{x}|\mathbf{z}_{0})$ be a Gaussian distribution with a small variance $\mathcal{N}(\mathbf{x};\mathbf{z}_{0},\delta^{2}I)$, where $\delta^{2}=10^{-4}$.
-This particular parameterization of the reverse process is not the only possibility.
-However, it is a convenient choice as it allows for the definition of the reverse process simply through the prediction of $\mathbf{x}$, akin to conventional diffusion models (Ho et al., 2020).
-3.3 Optimization
-We propose to optimize the forward and reverse processes of NFDM jointly, minimizing the following objective:
-$\mathcal{L}={\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}},$ (18)
-${\mathcal{L}_{\mathrm{rec}}}=\underset{q_{\varphi}(\mathbf{x},\mathbf{z}_{0})}{\mathbb{E}}\left[-\log p(\mathbf{x}|\mathbf{z}_{0})\right],$ (19)
-${\mathcal{L}_{\mathrm{diff}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right],$ (20)
-${\mathcal{L}_{\mathrm{prior}}}=\underset{q(\mathbf{x})}{\mathbb{E}}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})\|p(\mathbf{z}_{1})\big{)}\right].$ (21)
-As we demonstrate in Section A.1 this objectives provide a variational bound on the model’s likelihood $p_{\theta,\varphi}(\mathbf{x})$.
-The objective $\mathcal{L}$ shares similarities with the standard diffusion model objective (Ho et al., 2020), though the individual terms differ.
-Moreover, if the forward process is parameterized by $\varphi$, we need to optimize the objective (eq.
-18) with respect to these parameters.
-The objective $\mathcal{L}$ also exhibits strong connections with both the Flow Matching (Lipman et al., 2022) and Score Matching (Vincent, 2011) objectives.
-We explore these connections in Section B.2, where we also discuss the role of $g_{\varphi}$.
-A key characteristic of the NFDM objective is its amenability to training within a simulation-free paradigm, which is critical for efficient optimization.
-We summarize the training procedure in Algorithm 1.
+The above parameterization of the reverse process is not the only possibility.
+However, it is a convenient choice as it allows for the definition of the reverse process simply through the prediction of $\mathbf{x}$, similar to conventional diffusion models [19].
+We leave exploration of alternate parameterizations for future work.
+3.3 Optimization and Sampling
+With $F_{\varphi}$ in eq.
+7 parameterized such that $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})\approx p(\mathbf{z}_{1})$, we propose to optimize the forward and reverse processes of NFDM jointly, minimizing the following objective:
+$\mathcal{L}=\mathbb{E}_{u(t)q(\mathbf{x})q_{\varphi}(\mathbf{z}_{t}\mid\mathbf{x})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right].$ (12)
+Since the forward process is parameterized by $\varphi$, we need to optimize the objective (eq.
+12) with respect to both $\varphi$ and $\theta$ jointly.
+We discuss parameterization of the forward process in Section B.5.
+As we demonstrate in Section A.1 the objective $\mathcal{L}$ shares similarities with the standard diffusion model objective [19], and it provides a variational bound on the model’s log-likelihood $\log p_{\theta,\varphi}(\mathbf{x})$.
+The objective $\mathcal{L}$ also exhibits strong connections with both the Flow Matching [32] and Score Matching [67] objectives.
+We explore these connections in Section E.4, where we also discuss the role of $g_{\varphi}$ in detail.
+It is important to note that despite the parameterization of the reverse process through the prediction of $\mathbf{x}$, the objective $\mathcal{L}$ optimizes the generative dynamics and does not necessarily lead to accurate predictions of $\mathbf{x}$.
+A key characteristic of the NFDM objective is its compatibility with the simulation-free paradigm, which is critical for efficient optimization. We summarize the training procedure in Algorithm 1.
 Algorithm 1 Optimization of NFDM
-0:  $q(x)$, $F_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$
+$q(\mathbf{x})$, $F_{\varphi}$, $g_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$
 for learning iterations do
-     $\mathbf{x}\sim q(x)$, $t\sim u(t)$, $\{\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{1}\}\sim q(\cdot|\mathbf{x})$
-     ${\mathcal{L}_{\mathrm{rec}}}=-\log p(\mathbf{x}|\mathbf{z}_{0})$
-     ${\mathcal{L}_{\mathrm{diff}}}=\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}$
-     ${\mathcal{L}_{\mathrm{prior}}}=\log q_{\varphi}(\mathbf{z}_{1}|x)-\log p(\mathbf{z}_{1})$
-     Gradient step on $\theta$ and $\varphi$ w.r.t. ${\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}}$
+     $\mathbf{x}\sim q(\mathbf{x})$, $t\sim u(t)$
+     $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$
+     $\mathcal{L}=\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}$
+     Gradient step on $\theta$ and $\varphi$ w.r.t. $\mathcal{L}$
 end for
 Algorithm 2 Stochastic Sampling from NFDM
-0:  $F_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$, $T$ – number of steps
+$F_{\varphi}$, $g_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$, $T$ – number of steps
 $\Delta t=\frac{1}{T}$, $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$
@@
 $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$
-3.4 Sampling
-NFDM offers several methods for sampling from the trained reverse process.
-The first method is stochastic sampling, as defined in Section 3.2.
-This process involves the following steps: 1) sample $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$; 2) numerically simulate the SDE (eq.
-16) to obtain $\mathbf{z}_{0}$; 3) sample $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$.
+To sample from the trained reverse process we can simulate the SDE, as defined in Section 3.2.
 This procedure is summarized in Algorithm 2.
-Moreover, during the sampling process, we can adjust the level of stochasticity by modifying $g_{\varphi}(t)$ (eq.
-12).
+Additionally, during the sampling process, we can adjust the level of stochasticity by modifying $g_{\varphi}(t)$ (eq.
+9).
 It is important to note that changes to $g_{\varphi}(t)$ also influence $\tilde{f}^{F}_{\varphi}$ (eq.
-12) and $\hat{f}_{\theta,\varphi}$ (eq.
-17).
-In the extreme case where $g_{\varphi}(t)\equiv 0$, the reverse process becomes deterministic, allowing us to utilize off-the-shelf numerical ODE solvers and to estimate densities (Chen et al., 2018a; Grathwohl et al., 2018).
-3.5 Parameterization
-The parameterization of the variance function $g_{\varphi}(t)$ (eq.
-12) and the data point predictor $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ (eq.
-17) is straightforward.
-In our experiments, we parameterize them directly using neural networks.
-However, the parameterization of the transformation $F_{\varphi}$ (eq.
-9) is more complex.
-NFDM necessitates that $F_{\varphi}$ be differentiable and invertible with respect to $\varepsilon$, as well as provide access to the logarithm of the determinant of the transformation $\log|J_{F}^{-1}|$ (eq.
-14).
-Therefore, we use the following parameterization for $F_{\varphi}$:
-$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t),\quad\textrm{where}$ (22)
-$\mu_{\varphi}(\mathbf{x},t)$ $=(1-t)\mathbf{x}+t(1-t)\bar{\mu}_{\varphi}(\mathbf{x},t),$ (23)
-$\sigma_{\varphi}(\mathbf{x},t)$ $=\delta+t\bar{\sigma}_{\varphi}\big{(}(1-t)\mathbf{x},t\big{)}/\bar{\sigma}_{\varphi}(\mathbf{0},1).$ (24)
-In this context, $\bar{\mu}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ and $\bar{\sigma}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}_{+}$, with $\delta$ being a constant.
-In our experiments, we set $\delta^{2}=10^{-4}$.
-If $\bar{\mu}_{\varphi}$ and $\bar{\sigma}_{\varphi}$ are differentiable, this parameterization meets the requirements on $F_{\varphi}$ and renders $q(\mathbf{z}_{t}|\mathbf{x})$ a conditional Gaussian distribution with parameterized mean and diagonal covariance matrix.
-Furthermore, this reparameterization ensures that $q(\mathbf{z}_{0}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{0};\mathbf{x},\delta^{2}I)$ and $q(\mathbf{z}_{1}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{1};0,I)$.
-These restrictions are not necessary for $F_{\varphi}$, however, it eliminates the need to optimize the reconstruction loss ${\mathcal{L}_{\mathrm{rec}}}$ (eq.
-19) and the prior loss ${\mathcal{L}_{\mathrm{prior}}}$ (eq.
-21), as they are independent of the parameters $\varphi$ and $\theta$.
-The chosen parameterization of $F_{\varphi}$, $g_{\varphi}(t)$, and $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ is not the only possible approach and may not be optimal.
-For instance, it is possible to parameterize $F_{\varphi}$ as a flow-based model (Dinh et al., 2016; Kingma & Dhariwal, 2018), enabling a broader range of distributions than just conditional Gaussians.
-However, for the purpose of this paper, we maintain a simple setup and leave the exploration of the most effective parameterization for future research.
-We discuss the parameterization in more detail in Section C.2.
-Table 1: Comparison of NFDM results with baselines on density estimation tasks.
-We present results in terms of BPD, lower is better.
-NFDM achieves state-of-the-art results accross the three benchmark tasks.
-Model CIFAR10 ImageNet 32 ImageNet 64
-DDPM (Ho et al., 2020) $3.69$
-Score SDE (Song et al., 2020b) $2.99$
-Improved DDPM (Nichol & Dhariwal, 2021) $2.94$ $3.54$
-VDM (Kingma et al., 2021) $2.65$ $3.72$ $3.40$
-Score Flow (Song et al., 2021) $2.83$ $3.76$
-Flow Matching (Lipman et al., 2022) $2.99$ $3.53$ $3.31$
-Stochastic Interp. (Albergo & Vanden-Eijnden, 2022) $2.99$ $3.48$
-i-DODE (Zheng et al., 2023b) $2.56$ $3.43$
-MuLAN (Sahoo et al., 2023) $2.55$ $3.67$
-NFDM (this paper) $\mathbf{2.49}$ $\mathbf{3.36}$ $\mathbf{3.20}$
+9) and $\hat{f}_{\theta,\varphi}$ (eq.
+11).
+In the extreme case where $g_{\varphi}(t)\equiv 0$, the reverse process becomes deterministic, allowing us to utilize off-the-shelf numerical ODE solvers and to estimate densities [6, 16].
+We provide an extended sampling discussion in Section B.4.
+4 Neural Flow Bridge Models
+In this section we discuss a simple modification to the NFDM framework that enables us to learn bridges between two data distributions, $q(\mathbf{x}_{0})$ and $q(\mathbf{x}_{1})$, a modification we refer to as Neural Flow Bridge Models (NFBM).
+[…]
+$\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x}_{0},\mathbf{x}_{1}),$ (13)
+Similar to the discussion in Section 3.2, for the reverse process of NFBM, we predict the data points $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ from $\mathbf{z}_{t}$ and $t$, and substitute them into the conditional reverse time SDE (eq.
+10):
+$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}^{0,1}_{\theta}(\mathbf{z}_{t},t)\big{)}.$ (14)
+Here, the function $\hat{\mathbf{x}}^{0,1}_{\theta}(\mathbf{z}_{t},t)$ returns predictions of both $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$.
+Alternatively, the data point $\mathbf{x}_{1}$ can be reused as conditioning for intermediate steps of the reverse process instead of predicting both points, a strategy further detailed in Section C.2.
+When $F_{\varphi}$ in eq.
+13 is parameterized such that $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x}_{0},\mathbf{x}_{1})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})\approx p(\mathbf{z}_{1}|\mathbf{x}_{1})$ (see Section C.1 for parameterization details of the NFBM forward process), we propose training NFBM by minimizing the following objective:
+$\mathcal{L}=\mathbb{E}_{u(t)q(\mathbf{x}_{0},\mathbf{x}_{1})q_{\varphi}(\mathbf{z}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{1})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x}_{0},\mathbf{x}_{1})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right].$ (15)
+The derivation of this objective is provided in Section A.2.
+This objective shares key properties with the NFDM objective, it provides a variational bound on the model’s log-likelihood $\log p_{\theta,\varphi}(\mathbf{x}_{0})$, and is compatible with the simulation-free paradigm.
+Furthermore, it allows for sampling with various levels of stochasticity (by adjusting $g_{\varphi}(t)$) including deterministic sampling (when $g_{\varphi}(t)\equiv 0$).
+Thus, the NFDM framework not only enables the construction of generative models, but also facilitates learning bridge models between two data distributions.
 5 Restricted NFDM
-We introduce NFDM as a powerful framework that enables both pre-definition and learning of the forward process, potentially simplifying the task for the reverse process.
-However, there are in general an infinite number of forward and reverse processes that correspond to each other.
-So how to choose between them?
-The standard NFDM, as described in Section 3, learns one such pair of processes.
-However, the flexibility of NFDM opens up the possibility of choosing specific pairs of processes depending on the task.
-Suppose our objective is not just to learn any generative process for the data, but to learn one endowed with specific beneficial properties.
-For instance, one such property is straight generative ODE trajectories.
-Learning a generative process with straight ODE trajectories can be highly beneficial, as it enables the generation of samples with far fewer steps in the ODE solver.
-This can lead to significant computational savings.
-To achieve this, we must introduce some restrictions on the reverse process.
-One approach is to directly encode this in the definition of the reverse process.
-Alternatively, we can introduce penalties on the curvature of its trajectories.
-In both cases learnable forward process would then adapt to align with the reverse process.
+We introduce NFDM as a powerful framework that enables learning of the forward process.
+However, there is in general an infinite number of forward and reverse processes that correspond to each other.
+In this section we discuss how the flexibility of NFDM (and NFBM) allows learning generative dynamic with user-specified beneficial properties.
+Suppose our objective is to learn straight generative ODE trajectories, which can be highly beneficial, as it enables generation with far fewer steps in the ODE solver.
+One approach is to introduce penalties on the curvature of the reverse process’s trajectories and let the learnable forward process to adapt to align with the reverse process.
 As a result, we would obtain a generative process that not only corresponds to the forward process, ensuring accurate data generation, but also features the desired property of straight trajectories.
-We discuss NFDM with restrictions in more details in Section A.2.
-Importantly, conventional diffusion models are incapable of handling such constraints or penalization strategies.
-In diffusion models with a fixed forward process, the target for the reverse process is predetermined and the corresponding ODE trajectories are highly curved.
-Hence, imposing constraints on the reverse process, such as trajectory straightness, would lead to a mismatch with the forward process and, consequently, an inability to generate samples with high data fidelity.
-4.1 NFDM with optimal transport
-In this section, we discuss curvature penalization on the ODE trajectories of the reverse process in NFDM.
+We discuss NFDM with restrictions in more details in Section D.1.
+We propose learning a model with curvature penalty as suggested by [24]:
+${\mathcal{L}_{\mathrm{OT}}}=\mathcal{L}+\lambda{\mathcal{L}_{\mathrm{crv}}},\quad\textrm{where}\quad{\mathcal{L}_{\mathrm{crv}}}=\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left\|\frac{d\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{dt}\right\|_{2}^{2}.$ (16)
 We refer to this variant as NFDM-OT [^1] OT stands for optimal transport.
@@
 While straight trajectories are a necessary condition for dynamic optimal transport, they are not sufficient..
-We propose learning a model with an additional curvxature penalty as suggested by Kelly et al. (2020):
-${\mathcal{L}_{\mathrm{OT}}}$ $=\mathcal{L}+\lambda{\mathcal{L}_{\mathrm{crv}}},\quad\textrm{where}$ (25)
-${\mathcal{L}_{\mathrm{crv}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}$ $\left\|\frac{d\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{dt}\right\|_{2}^{2}.$ (26)
 ${\mathcal{L}_{\mathrm{crv}}}$ is an additional curvature loss that penalizes the second time derivative of the generative ODE trajectories.
@@
 ${\mathcal{L}_{\mathrm{crv}}}\equiv 0$ ensures that the generative trajectories are straight lines.
-Although we parameterize $\hat{f}_{\theta,\varphi}$ (eq.
-17) through $\tilde{f}^{B}_{\varphi}$ (eq.
-15), ${\mathcal{L}_{\mathrm{crv}}}$ does not penalize the curvature of the conditional forward trajectories.
-The penalty requires $F_{\varphi}$ to be twice differentiable with respect to time.
-The specifics of the curvature loss are elaborated upon in more detail in Section A.2.
-Even though the parameterization, as discussed in section 3.5, is simple and does not guarantee reaching the zero curvature, we find that in practice it is flexible enough to successfully learn straight line generative trajectories.
 In our experiments, we set $\lambda=10^{-2}$.
 Empirical evidence suggests that higher values of $\lambda$ lead to slower convergence of the model.
-We provide empirical results for NFDM-OT in Section 5.2.
+The specifics of the curvature loss are elaborated upon in more detail in Section D.2.
+We would like to emphasise that the purpose of this section is to provide an example of how the NFDM framework allows learning dynamics with specific properties.
+Not to propose the best way for learning generative straight-line generative dynamics.
+In addition, we note that conventional diffusion models are incapable of handling such penalization strategies.
+In diffusion models with a fixed forward process, the target for the reverse process is predetermined and the corresponding ODE trajectories are highly curved.
+Hence, imposing constraints on the reverse process, such as trajectory straightness, would lead to a mismatch with the forward process and, consequently, an inability to generate samples with high data fidelity.
+6 Experiments
+We first showcase results demonstrating that NFDM consistently achieves better likelihood compared to baselines, obtaining state-of-the-art diffusion modeling results on the CIFAR-10 [30] and downsampled ImageNet [11, 66] datasets.
+Then, we explore the NFDM-OT modification, which penalizes the curvature of the deterministic generative trajectories.
+The NFDM-OT reduces trajectory curvature, significantly lowering the number of generative steps required for sampling.
+Finally, we demonstrate the NFBM modification, learning bridges on the AFHQ [8] dataset and several synthetic examples.
+We report the NLL in Bits Per Dimension (BPD) and sample quality measured by the Frechet Inception Distance (FID) score [18].
+The NLL is calculated by integrating the ODEs using the RK45 solver [15], with all NLL metrics computed on test data.
+For FID, we provide the average over 50k generated images.
+Unless otherwise stated, we parameterized the forward process of NFDM and NFBM such that $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is a Gaussian with learnable mean and covariance (see details in Sections B.5 and C.1).
+For a detailed description of parameterizations and other experimental details, please refer to Appendix F.
+The primary aim of the experiments with NFDM-OT and NFBM is not to introduce a novel model that surpasses others in few-step generation or learning bridges, but rather to showcase the ability of the NFDM framework to learn generative dynamics with specific properties.
+The straightness of the trajectories is just one example of such a property.
+We leave it for future research to explore different parameterizations and modified objectives for NFDM that may yield even better results.
+Please refer to section G.3 for generated samples.
+6.1 Likelihood Estimation
+Table 1: Comparison of NFDM results with baselines on density estimation tasks.
+We present results in terms of BPD, lower is better.
+NFDM achieves state-of-the-art results across all three benchmark tasks.
+Model CIFAR10 ImageNet 32 ImageNet 64
+DDPM [19] $3.69$
+Score SDE [60] $2.99$
+Improved DDPM [38] $2.94$ $3.54$
+VDM [27] $2.65$ $3.72$ $3.40$
+Score Flow [58] $2.83$ $3.76$
+Flow Matching [32] $2.99$ $3.53$ $3.31$
+Stochastic Interp. [2] $2.99$ $3.48$
+i-DODE [74] $2.56$ $3.43$
+NDM [4] $2.70$ $3.55$ $3.35$
+MuLAN [48] $2.55$ $3.67$
+NFDM (Gaussian $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$) $2.49$ $3.36$ $\mathbf{3.20}$
+NFDM (non-Gaussian $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$) $\mathbf{2.48}$ $\mathbf{3.34}$ $\mathbf{3.20}$
+(a) Score SDE [60]
+(b) NFDM-OT (this paper)
+Figure 1: Comparison of trajectories between the data distribution (on the left) and the prior distribution (on the right), as learned by conventional diffusion and NFDM-OT.
+For the first experiments, in addition to Gaussian we also provide results with non-Gaussian parameterization of the forward process (see Section B.5).
+Table 1 summarizes the NLL results on the CIFAR-10 and two downsampled ImageNet datasets.
+Notably, NFDM outperforms diffusion-based baselines on all three datasets with both parameterizations, achieving state-of-the-art performance.
+The improved performance due to NFDM is a natural progression for the following reasons.
+First, it is well-established [58, 74] that diffusion models exhibit improved likelihood estimation when trained with the full ELBO objective.
+The objective $\mathcal{L}$ (eq.
+12) used for training the NFDM is also a variational bound on the likelihood.
+Second, diffusion models can be seen as hierarchical VAEs.
+From this perspective, most baselines in Table 1 resemble VAEs with either fixed or constrained variational distributions.
+In contrast, the NFDM extends beyond these baselines by providing a more flexible variational distribution.
+That is true even for the Gaussian parameterization with a mean and covariance that is non-linear in $\mathbf{x}$ and $t$.
+This flexibility allows the NFDM to better conform to the reverse process, consequently enhancing likelihood estimation.
+6.2 Straight Trajectories
+We next evaluate NFDM-OT, designed to penalize the deterministic generative trajectory curvature.
+First, we compare NFDM-OT with a conventional continuous-time diffusion model.
+Figure 1(a) illustrates deterministic trajectories between a two-dimensional data distribution and a unit Gaussian distribution learnt by a conventional diffusion model [60].
+Figure 1(b) depicts trajectories learnt by NFDM-OT.
+Conventional diffusion, being constrained in its forward process, learns highly curved trajectories, whereas NFDM-OT successfully learns straight generative trajectories as desired.
+In Section G.1 we provide some additional reult, demonstrating the importance of learnable forward process for NFDM-OT.
+Then, we present results of NFDM-OT on image datasets.
+Table 2 reports the FID scores for $2$, $4$, and $12$ Number of Function Evaluations (NFE) with respect to the function $\hat{f}_{\theta,\varphi}$ (eq.
+11).
+In this experiment, we employ Euler’s method for sampling integration.
+For the specified NFEs, NFDM-OT demonstrates superior sample quality compared to other approaches with similar NFE values.
+Specifically, NFDM-OT outperforms approaches specifically designed to minimize the curvature of generative trajectories [43, 31].
+Importantly, NFDM-OT is trained with an ELBO-based objective (eq.
+12), which is known to yield higher FID scores for diffusion models [58, 74].
+In contrast, some of the approaches listed in Table 2 are trained with different objectives, leading to improved FID scores.
+Even so, for comparable NFE values NFDM-OT still achieves superior results.
+We provide additional results of NFDM-OT in Section G.2.
 Table 2: Summary of FID results for few-step generation.
@@
 Model NFE $\downarrow$ FID $\downarrow$ NFE $\downarrow$ FID $\downarrow$ NFE $\downarrow$ FID $\downarrow$
-DDPM (${\mathcal{L}_{\mathrm{simple}}}$) (Ho et al., 2020) $1000$ $3.17$
-DDPM (ELBO) (Ho et al., 2020) $1000$ $13.51$
-Flow Matching (Lipman et al., 2022) $142$ $6.35$ $122$ $5.02$ $138$ $14.14$
-DDIM (Song et al., 2020a) $10$ $13.36$
+DDPM (${\mathcal{L}_{\mathrm{simple}}}$) [19] $1000$ $3.17$
+DDPM (ELBO) [19] $1000$ $13.51$
+Flow Matching [32] $142$ $6.35$ $122$ $5.02$ $138$ $14.14$
+DDIM [56] $10$ $13.36$
 $12$ $5.28$
-DPM Solver (Lu et al., 2022) $24$ $2.75$
-Trajectory Curvature Minimization (Lee et al., 2023) $5$ $18.74$
+DPM Solver [37] $24$ $2.75$
+Trajectory Curvature Minimization [31] $5$ $18.74$
 $4$ $17.28$ $4$ $38.45$
-Multisample Flow Matching (Pooladian et al., 2023) $12$ $7.18$ $12$ $17.6$
+Multisample Flow Matching [43] $12$ $7.18$ $12$ $17.6$
 $2$ $12.44$ $2$ $9.83$ $2$ $27.70$
@@
 NFDM-OT (this paper) $12$ $5.20$ $12$ $4.11$ $12$ $11.58$
-5 Experiments
-We first showcase results demonstrating that NFDM consistently achieves better likelihood compared to baselines, obtaining state-of-the-art diffusion modeling results on the CIFAR-10 and downsampled ImageNet datasets.
-Then, we explore the NFDM-OT modification, which penalizes the curvature of the deterministic generative trajectories.
-The NFDM-OT reduces trajectory curvature, significantly lowering the number of generative steps required for sampling.
-We evaluate NFDM on synthetic data, CIFAR-10 (Krizhevsky et al., 2009), and two downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) datasets.
-For parameterization we use multi-layer perceptrons for synthetic data and the U-Net architecture from Dhariwal & Nichol (2021) for images.
-We report the NLL in Bits Per Dimension (BPD) and sample quality measured by the FID score (Heusel et al., 2017).
-The NLL is calculated by integrating the ODEs using the RK45 solver Dormand & Prince (1980), with all NLL metrics computed on test data.
-For the FID, we provide the average over 50k images.
-For a detailed description of parameterizations and other experimental details, please refer to Appendix C.
-(a) Score SDE (Song et al., 2020b)
-(b) NFDM (this paper)
-Figure 1: Comparison of trajectories between the data distribution (on the left) and the prior distribution (on the right).
-5.1 Likelihood estimation
-We have trained NFDM on the CIFAR-10 and downsampled ImageNet datasets.
-The NLL results are summarized in Table 1.
-Notably, NFDM outperforms diffusion-based baselines on all three datasets.
-We consider this enhanced performance of the NFDM a natural progression due to the following reasons.
-First, it is well-established (Song et al., 2021; Zheng et al., 2023b) that diffusion models exhibit improved likelihood estimation when trained with the full ELBO objective.
-As delineated in Section 3.3, the objective $\mathcal{L}$ (eq.
-18) used for training the NFDM constitutes a variational bound on the likelihood.
-Second, as previously mentioned, diffusion models can be seen as hierarchical VAEs.
-From this perspective, the baselines in Table 1 resemble VAEs with either fixed or constrained variational distributions.
-In contrast, the NFDM extends beyond these baselines by providing a more flexible variational distribution.
-This flexibility allows the NFDM to better conform to the reverse process, consequently enhancing likelihood estimation.
-5.2 Straight trajectories
-We next evaluate NFDM-OT, which is designed to penalize the curvature of deterministic generative trajectories.
-Initially, we compare NFDM-OT with a conventional continuous-time diffusion model.
-Figure 1(a) illustrates deterministic trajectories between a two-dimensional data distribution and a unit Gaussian distribution learnt by a conventional diffusion model (Song et al., 2020b).
-Figure 1(b) depicts trajectories learnt by NFDM-OT.
-Conventional diffusion, being constrained in its forward process, learns highly curved trajectories, whereas NFDM-OT successfully learns straight generative trajectories as desired.
-Subsequently, we present evaluation results of NFDM-OT on image datasets.
-Table 2 reports the FID scores for $2$, $4$, and $12$ Number of Function Evaluations (NFE) with respect to the function $\hat{f}_{\theta,\varphi}$ (eq.
-16).
-In this experiment, we employ Euler’s method for sampling integration.
-For the specified NFEs, NFDM-OT demonstrates superior sample quality compared to other approaches with similar NFE values.
-Specifically, NFDM-OT outperforms approaches that minimize the curvature of generative trajectories (Pooladian et al., 2023; Lee et al., 2023).
-Importantly, NFDM-OT is trained with an ELBO-based objective (eq.
-18), which is known to yield higher FID scores for diffusion models (Song et al., 2021; Zheng et al., 2023b).
-In contrast, some of the approaches listed in Table 2 are trained with different objectives, leading to improved FID scores.
-Even so, for comparable NFE values NFDM-OT still achieves superior results.
-The primary aim of the experiments with NFDM-OT is not to introduce a novel model that surpasses others in few-step generation, but rather to showcase NFDM’s ability to learn generative dynamics with specific properties.
-The straightness of the trajectories is just one example of such properties.
-We leave it for future research to explore different parameterizations and a modified objectives for NFDM that may yield even better results.
+Figure 2: Trajectories of NFBM between two-dimensional data distributions, trained with an additional penalty to avoid obstacles.
+6.3 Bridges
+Table 3: FID values of dog $\rightarrow$ cat on AFHQ $64$.
+Model FID $\downarrow$
+DSBM [53] $14.16$
+GSBM [33] $12.39$
+NFBM $\mathbf{12.06}$
+We consider the task of translating images of dogs into images of cats from the downsampled AFHQ dataset.
+Table 3 reports the FID scores on images sampled with $1000$ steps.
+Notably, NFBM outperforms the baselines, demonstrating the effectiveness of our framework for learning bridges.
+Finally, we demonstrate that the framework flexibility also allows for learning dynamics with specific properties in the case of NFBM.
+To illustrate this, we adopted the strategy discussed in Section 5 and learn NFBM with an additional penalty to avoid obstacles (see details in Section F.2).
+Figure 2 visualizes the learned stochastic trajectories.
+It is evident that NFBM has efficiently learned to translate distributions while avoiding obstacles.
 7 Related Work
-Diffusion models, originally proposed by Sohl-Dickstein et al.
-(2015), have evolved significantly through subsequent developments (Song & Ermon, 2019; Ho et al., 2020).
-These advancements have resulted in remarkable generative quality for high-dimensional data distributions (Dhariwal & Nichol, 2021; Saharia et al., 2022).
+Diffusion models, originally proposed by [55], have evolved significantly through subsequent developments [59, 19].
+These advancements have resulted in remarkable generative quality for high-dimensional data distributions [12, 47].
 Nevertheless, conventional diffusion models, typically relying on a linear Gaussian forward process, may not optimally fit some data distributions.
-To address this, alternative forward processes have been explored, such as combining blurring with Gaussian noise injection (Rissanen et al., 2022; Daras et al., 2022; Hoogeboom & Salimans, 2022), diffusion in the wavelet spectrum (Phung et al., 2023), and forward processes based on the exponential family (Okhotin et al., 2023).
+To address this, alternative forward processes have been explored, such as combining blurring with Gaussian noise injection [45, 9, 22], diffusion in the wavelet spectrum [42], and forward processes based on the exponential family [40].
 These models are limited by their fixed forward processes and may be seen as specific instances of NFDM.
 Some studies have focused on making the forward process learnable.
-Approaches include a learnable noise injection schedule (Kingma et al., 2021; Nichol & Dhariwal, 2021; Sahoo et al., 2023) and learning data transformations like time-independent transformations based on VAEs (Vahdat et al., 2021; Rombach et al., 2022) and normalizing flows (Kim et al., 2022), or time-dependent transformations (Gu et al., 2022; Nielsen et al., 2023; Bartosh et al., 2023).
+Approaches include a learnable noise injection schedule [27, 38, 48] and learning data transformations like time-independent transformations based on VAEs [65, 46] and normalizing flows [25], or time-dependent transformations [17, 54, 39, 4].
 These methods can also be considered special cases of NFDM with specific transformations $F_{\varphi}$ (eq.
-9).
-Recent studies (Zhang & Chen, 2021; De Bortoli et al., 2021; Shi et al., 2023) have explored generative models based on Schrödinger Bridge theory and finite-time diffusion constructions.
-While these models offer novel approaches to learning forward transformations, in contrast to NFDM, they depart from the simulation-free paradigm by requiring the full simulation of the stochastic processes for training, increasing their computational cost.
-Despite the simulation-free nature of the training procedure, diffusion models still necessitate full reverse process simulations for sample generation, leading to slow and computationally expensive inference.
-To address this, in an orthogonal line of works alternative sampling methods have been studied, such as deterministic sampling in discrete (Song et al., 2020a) and continuous time (Song et al., 2020b) and novel numerical solvers (Tachibana et al., 2021; Liu et al., 2022a; Lu et al., 2022; Shaul et al., 2023b).
-Additionally, distillation techniques have been applied to both discrete (Salimans & Ho, 2022) and continuous time models (Song et al., 2023; Zheng et al., 2023a; Yin et al., 2023) to enhance sampling speed, albeit at the cost of training additional models.
+7).
+To improve the sampling efficiency, in orthogonal line of works alternative sampling methods have been studied [61, 34, 37, 52].
+Additionally, distillation techniques have been applied [49, 57, 73, 71] to enhance sampling speed, albeit at the cost of training additional models.
 Most of these approches are compatible with NFDM and may be combined for further gains.
-The significance of deterministic trajectories in efficient sampling was highlighted by Karras et al.
-(2022).
-At the same time, building on diffusion model concepts, Lipman et al.
-(2022); Liu et al.
-(2022b); Albergo & Vanden-Eijnden (2022) introduced simulation-free methods for learning deterministic generative dynamics.
-Based on these ideas, Liu et al.
-(2022b) and Liu (2022) proposed distillation procedures to straighten deterministic generative trajectories.
-Since these methods are based on distillation, we consider them orthogonal to NFDM.
-Pooladian et al.
-(2023) and Tong et al.
-(2023) proposed to construct the forward process with optimal data-noise couplings to learn straighter generative trajectories.
-While this method is theoretically justified, it relies on estimating an optimal coupling over minibatches of the entire dataset, which, for large datasets, may become uninformative as to the true coupling.
-In contrast, NFDM-OT directly penalizes curvature of generative trajectories.
-Lee et al.
-(2023) suggested learning the forward process distribution $q(\mathbf{z}_{1}|\mathbf{x})$ and interpolating linearly between $\mathbf{x}$ and $\mathbf{z}_{1}$.
-Similarly, Shaul et al.
-(2023a) aimed to modify the forward process to reduce curvature in the generative process.
-Both approaches can be viewed as specific instances of NFDM.
-Lastly, Albergo et al.
-(2023) suggested optimizing the forward process simultaneously to training the reverse process, resultin in a challenging high dimentional min-max problem.
-Moreover, only the theoretical algorithm was presented, without empirical study and descriptions of ways to parameterize such a learnable forward process.
-We continue discussing related works in Appendix B.
-7 Discussion and Limitations
+Another line of works proposed technics for learning straighter generative trajectories.
+However, these approaches requires distillation [36, 35], solving unstable min-max problem [1], estimating an optimal coupling over minibatches [43, 63] of the entire dataset, which, for large datasets, may become uninformative, or can be viewed as specific instances [31, 51] of NFDM.
+The connections between NFDM and some related works is discussed further in Appendix E.
+8 Conclusion and Limitations
 In this paper we introduced NFDM, a novel simulation-free framework for improved diffusion modeling through a learnable forward processes.
-Our approach outperforms baseline diffusion models on standard benchmarks, showcasing the effectiveness of NFDM in log-likelihood estimation.
-We introduce NFDM not merely as a specific model, but rather as a versatile framework that facilitates the predefined specification and learning of the forward process.
-For the sake of simplicity in this work, we have chosen a straightforward Gaussian parameterization for the forward process (see Section 3.5).
-However, NFDM is capable of supporting far more flexible parameterizations beyond standard Gaussian distributions.
-Similarly, the penalty on the curvature of deterministic trajectories discussed in Section 4.1 represents just one specific example of the restrictions that NFDM can accommodate.
+NFDM outperforms baseline diffusion models on standard benchmarks, showcasing its effectiveness.
+However, NFDM is not merely just another model, but rather it is a versatile framework that facilitates the predefined specification and learning of a forward process.
+Similarly, the penalty on the curvature of deterministic trajectories discussed in Section 5 represents just one specific example of the restrictions that NFDM can accommodate.
+Nonetheless, the advantages of NFDM come with certain trade-offs.
+Once the forward process is parameterized using a neural network, this leads to approximately 2.2 times longer optimization iteration of NFDM compared to conventional diffusion models.
+Additionally, as discussed in Section 3.1, NFDM doesn’t allow parameterisation of the forward process with functions, that doesen’t match certain conditions.
+Despite these challenges, we are optimistic about the potential of NFDM.
 NFDM opens up significant prospects for exploring new generative dynamics.
-Nonetheless, the advantages of NFDM come with certain trade-offs.
-Once the forward process is parameterized using a neural network, this leads to increased computational costs compared to conventional diffusion models.
-In our experiments, an optimization iteration of NFDM takes approximately 2.2 times longer than that of conventional diffusion models.
-Despite these challenges, we are optimistic about the potential of NFDM in various domains and practical applications, given its flexibility in learning generative processes.
-We also believe that with alternative parameterizations, modifications of the objective, and the integration of orthogonal approaches like distillation, NFDM has the potential to achieve even better results.
+We believe that with alternative parameterizations, modifications of the objective, and the integration of orthogonal approaches, NFDM has the potential to achieve even better results across a range of data modalities.
 Appendix A Derivations
-A.1 Derivation of objective
-For the forward process, we have the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which is implicitly defined by the invertible transformation $\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x})$ (eq.
-9), where $\varepsilon\sim q(\varepsilon)$.
-Additionally, there’s a SDE (eq.
-15) starting from $\mathbf{z}_{1}\sim q(\mathbf{z}_{1}|\mathbf{x})$ and flowing backwards in time:
-$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w}.$ (27)
-This SDE corresponds to the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
-The reverse process involves a prior distribution $p(\mathbf{z}_{1})$ and a reverse SDE (eq. 15):
+A.1 Derivation of NFDM’s Objective
+Our goal is to derive a variational bound on the negative log-likelihood of NFDM.
+For the convenience of the following derivations we will consider the forward process as a conditional reverse SDE (eq.
+10) that starts from $\mathbf{z}_{1}\sim q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ and flows backwards in time:
+$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\bar{\mathbf{w}}.$ (17)
+Similarly to a conditional SDE (eq.
+9) flowing forwards in time, this SDE also corresponds to conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which is implicitly defined by the invertible transformation $\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x})$ (eq.
+7), where $\varepsilon\sim q(\varepsilon)$.
+For the reverse process we use the definition from Section 3.2 as reverse SDE (eq.
+11) that starts from the prior distribution $p(\mathbf{z}_{1})$, flows backwards in time, and finally samples $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$:
 $d\mathbf{z}_{t}=$ $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}}.$ (18)
-Our goal is to derive a variational bound on the negative log-likelihood of the model.
-We begin by discretizing the conditional reverse SDE in eq.
+To derive the NFDM’s variational bound on the NLL we discretize processes, derive an objective in discrete case and then consider the limit of discretization to switch back to continuous time.
+We discretize the conditional reverse SDE in eq.
 17 and the reverse SDE in eq.
-28, transitioning from continuous-time trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ to discrete-time trajectories $\bar{\mathbf{z}}_{0},\bar{\mathbf{z}}_{\frac{1}{T}},\dots,\bar{\mathbf{z}}_{1}$, where $T$ represents the number of discrete steps.
-For convenience, we also introduce $\Delta t=\frac{1}{T}$.
-The discretization of the conditional reverse SDE is as follows:
-$\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})=q_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x}),\quad\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x}),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (29)
-We similarly discretize the reverse SDE:
-$\bar{p}(\bar{\mathbf{z}}_{1})=p(\bar{\mathbf{z}}_{1}),\quad\bar{p}(\mathbf{x}|\bar{\mathbf{z}}_{0})=p(\mathbf{x}|\bar{\mathbf{z}}_{0}),\quad\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (30)
+18, transitioning from continuous-time trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ to discrete-time trajectories $\bar{\mathbf{z}}_{0},\bar{\mathbf{z}}_{\frac{1}{T}},\dots,\bar{\mathbf{z}}_{1}$, where $T$ represents the number of discrete steps of size $\Delta t=\frac{1}{T}$.
+The discrete conditional reverse process starts from distribution $\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})=q_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})$ and follows conditional reverse Markov chain:
+$\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x}),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (19)
+Similarly, the discrete generative process shares distributions $\bar{p}(\bar{\mathbf{z}}_{1})=p(\bar{\mathbf{z}}_{1})$ and $\bar{p}(\mathbf{x}|\bar{\mathbf{z}}_{0})=p(\mathbf{x}|\bar{\mathbf{z}}_{0})$ with the continuous one.
+However, instead of following the reverse SDE (eq.
+18) it follows the reverse Markov chain:
+$\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (20)
 It’s important to note that once we discretize the trajectories, $\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t}|\mathbf{x})\neq q_{\varphi}(\bar{\mathbf{z}}_{t}|\mathbf{x})$ and $\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t})\neq p_{\theta,\varphi}(\bar{\mathbf{z}}_{t})$ except for $t=1$.
@@
 Thus, we know that as $\Delta t\rightarrow 0$, the discretized trajectories converge to the continuous ones.
-Now, considering these discrete processes and recognizing that they are Markovian conditional and unconditional processes moving backwards in time, we can leverage a key result from Ho et al. (2020):
-$\mathbb{E}_{q(\mathbf{x})}\left[-\log\bar{p}_{\theta,\varphi}(\mathbf{x})\right]\leq\bar{\mathcal{L}}_{\mathrm{recon}}+\bar{\mathcal{L}}_{\mathrm{diff}}+\bar{\mathcal{L}}_{\mathrm{prior}},\quad\textrm{where}$ (31)
-¯$\bar{\mathcal{L}}_{\mathrm{recon}}=\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{0})}\left[-\log p(\mathbf{x}|\bar{\mathbf{z}}_{0})\right],$ (32)
+Now, considering these discrete processes and recognizing that they are Markovian conditional and unconditional processes moving backwards in time, we can leverage a key result from [55, 19]:
+$\mathbb{E}_{q(\mathbf{x})}\left[-\log\bar{p}_{\theta,\varphi}(\mathbf{x})\right]\leq{\bar{\mathcal{L}}_{\mathrm{rec}}}+{\bar{\mathcal{L}}_{\mathrm{diff}}}+{\bar{\mathcal{L}}_{\mathrm{prior}}},\quad\textrm{where}$ (21)
+¯${\bar{\mathcal{L}}_{\mathrm{rec}}}=\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{0})}\left[-\log p(\mathbf{x}|\bar{\mathbf{z}}_{0})\right],$ (22)
 ¯${\bar{\mathcal{L}}_{\mathrm{diff}}}=\sum_{t=\frac{1}{T}}^{T}\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{t})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})\|\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})\big{)}\right],$ (23)
 ¯${\bar{\mathcal{L}}_{\mathrm{prior}}}=\mathbb{E}_{q(\mathbf{x})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})\|\bar{p}(\bar{\mathbf{z}}_{1})\big{)}\right].$ (24)
-Consequently, eq.
-31 establishes a variational upper bound on the negative log-likelihood of discretized reverse process.
-Taking the limits of this upper bound, the reconstruction term ¯$\bar{\mathcal{L}}\mathrm{recon}$ and the prior term ¯$\bar{\mathcal{L}}\mathrm{prior}$ retain their forms.
-However, we can reformulate the diffusion term:
+Equation 21 establishes a variational upper bound on the negative log-likelihood of discretized reverse process.
+Taking the limits of this upper bound, the reconstruction term ${\mathcal{L}_{\mathrm{rec}}}={\bar{\mathcal{L}}_{\mathrm{rec}}}$ and the prior term ${\mathcal{L}_{\mathrm{prior}}}={\bar{\mathcal{L}}_{\mathrm{prior}}}$ retain their forms.
+However, we can rewrite the diffusion term:
 $\mathcal{L}_{\mathrm{diff}}$ $=\lim_{\Delta t\rightarrow 0}{\bar{\mathcal{L}}_{\mathrm{diff}}}$ (25)
@@
 where $u(t)$ is a uniform distribution over a unit interval.
-Therefore, as the discretized trajectory tends to the continuous one when $\Delta t\rightarrow 0$, we can use the obtained limits as a variational bound on the negative log-likelihood of the continuous model.
-A.2 NFDM with restrictions
+Therefore, as the discretized trajectory tends to the continuous one when $\Delta t\rightarrow 0$, we can use the obtained limits as a variational bound on the negative log-likelihood of the continuous model:
+$\mathbb{E}_{q(\mathbf{x})}\left[-\log p_{\theta,\varphi}(\mathbf{x})\right]\leq\mathcal{L}={\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}}.$ (30)
+In general case, when function $F_{\varphi}$ is actually parameterized by $\varphi$, the objective $\mathcal{L}$ implies that we must optimise all three terms with respect to parameters $\varphi$ and $\theta$.
+However, when $F_{\varphi}$ is parameterized such that $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})\approx p(\mathbf{z}_{1})$ for any $\varphi$, the reconstruction ${\mathcal{L}_{\mathrm{rec}}}$ and prior ${\mathcal{L}_{\mathrm{prior}}}$ terms are constants with respect to the parameters.
+A.2 Derivation of NFBM’s Objective
+Similarly to the derivation of the NFDM objective in Section A.1, here we also apply discretization to derive the objective of NFBM.
+The conditional reverse process (eq.
+17) and it’s discretization (eq.
+19) stays the same, except now the conditional reverse process are conditioned on a pair of points $\mathbf{x}=(\mathbf{x}_{0},\mathbf{x}_{1})$, sampled from a joined data distribution $(\mathbf{x}_{0},\mathbf{x}_{1})\sim q(\mathbf{x}_{0},\mathbf{x}_{1})$.
+The reverse process (eq.
+18) and its discretization (eq.
+20) also have the same structure.
+However, for NFBM the reverse process starts from distribution $\bar{p}(\mathbf{z}_{1}|\mathbf{x}_{1})=p(\mathbf{z}_{1}|\mathbf{x}_{1})$ and at the end samples from $\bar{p}(\mathbf{x}_{0}|\mathbf{z}_{0})=p(\mathbf{x}_{0}|\mathbf{z}_{0})$.
+We derive the variational bound on NLL of NFBM:
+$\mathbb{E}$ [${}_{q(\mathbf{x}_{0})}\left[-\log\bar{p}_{\theta,\varphi}(\mathbf{x}_{0})\right]$ (31)
+[…]
+$=$ $\mathbb{E}_{q(\mathbf{x}_{0})}\left[-\log\mathbb{E}_{\bar{q}(\mathbf{x}_{1},\bar{\mathbf{z}}_{0:1}|\mathbf{x}_{0})}\left[\bar{p}(\mathbf{x}_{0}|\bar{\mathbf{z}}_{0})\frac{\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{0:1-\Delta t}|\bar{\mathbf{z}}_{1})\bar{p}(\bar{\mathbf{z}}_{1}|\mathbf{x}_{1})q(\mathbf{x}_{1})}{\bar{q}_{\varphi}(\bar{\mathbf{z}}_{0:1-\Delta t}|\bar{\mathbf{z}}_{1},\mathbf{x}_{0},\mathbf{x}_{1})\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})q(\mathbf{x}_{1}|\mathbf{x}_{0})}\right]\right]$ (33)
+$\leq$ $\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{z}_{0:1})}\left[-\log\bar{p}(\mathbf{x}_{0}|\bar{\mathbf{z}}_{0})\frac{\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{0:1-\Delta t}|\bar{\mathbf{z}}_{1})\bar{p}(\bar{\mathbf{z}}_{1}|\mathbf{x}_{1})q(\mathbf{x}_{1})}{\bar{q}_{\varphi}(\bar{\mathbf{z}}_{0:1-\Delta t}|\bar{\mathbf{z}}_{1},\mathbf{x}_{0},\mathbf{x}_{1})\bar{q}(\bar{\mathbf{z}}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})q(\mathbf{x}_{1}|\mathbf{x}_{0})}\right]$ (34)
+$=$ ¯${\bar{\mathcal{L}}_{\mathrm{rec}}}+{\bar{\mathcal{L}}_{\mathrm{diff}}}+{\bar{\mathcal{L}}_{\mathrm{prior}}}+\textrm{const},\quad\textrm{where}$ (35)
+¯${\bar{\mathcal{L}}_{\mathrm{rec}}}=\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{z}_{0})}\left[-\log\bar{p}(\mathbf{x}_{0}|\bar{\mathbf{z}}_{0})\right],$ (36)
+¯${\bar{\mathcal{L}}_{\mathrm{diff}}}=\sum_{t=\frac{1}{T}}^{T}\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},\bar{\mathbf{z}}_{t})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x}_{0},\mathbf{x}_{1})\|\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})\big{)}\right],$ (37)
+¯${\bar{\mathcal{L}}_{\mathrm{prior}}}=\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})\|\bar{p}(\bar{\mathbf{z}}_{1}|\mathbf{x}_{1})\big{)}\right],$ (38)
+$\textrm{const}=\mathbb{E}_{q(\mathbf{x}_{0})}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}q(\mathbf{x}_{1}|\mathbf{x}_{0})\|q(\mathbf{x}_{1})\big{)}\right].$ (39)
+Here, the reconstruction term ${\mathcal{L}_{\mathrm{rec}}}={\bar{\mathcal{L}}_{\mathrm{rec}}}$ and the prior term ${\mathcal{L}_{\mathrm{prior}}}={\bar{\mathcal{L}}_{\mathrm{prior}}}$ do not depend on discretization, and the constant term does not depend on model’s parameters.
+For the diffusion term ¯${\bar{\mathcal{L}}_{\mathrm{diff}}}$ it is easy to see that in the case of NFBM it has the same structure as in NFDM.
+Therefore, taking the limit of the number of discretization steps $T$ the diffusion term ¯${\bar{\mathcal{L}}_{\mathrm{diff}}}$ transforms to:
+$\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{u(t)\bar{q}_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},\bar{\mathbf{z}}_{t})}\left[\frac{1}{2g^{2}_{\varphi}(t)}\left\|\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x}_{0},\mathbf{x}_{1})-\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t)\right\|_{2}^{2}\right],$ (40)
+where $u(t)$ is a uniform distribution over a unit interval.
+This gives us a variational bound on the negative log-likelihood of the NFBM, with a very similar structure as the NFDM objective.
+Additionally, for NFBM we also may omit optimization of ${\mathcal{L}_{\mathrm{rec}}}$ and prior ${\mathcal{L}_{\mathrm{prior}}}$ when $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x}_{0},\mathbf{x}_{1})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})\approx p(\mathbf{z}_{1}|\mathbf{x}_{1})$.
+Appendix B Details of NFDM
+B.1 Motivation of Generalized Forward Process
+In this section, we provide an extended discussion on the motivation behind the learnable forward process.
+From a theoretical perspective, diffusion models can be viewed as a hierarchical VAE, where the objective function of diffusion models is the variational bound on model log-likelihood.
+Therefore, diffusion models aim to match the joint distributions of data $\mathbf{x}$ and latent variables $\mathbf{z}_{0:1}$ (for simplicity, let’s consider a discrete-time model with some $T$ time steps).
+However, as we know from [29, 44], the variational objective is equal to:
+$\underbrace{{\mathrm{D}_{\mathrm{KL}}}\big{(}q_{\varphi}(\mathbf{x},\mathbf{z}_{0:1})\|p_{\theta}(\mathbf{x},\mathbf{z}_{0:1})\big{)}}_{\textrm{objective}}=\underbrace{{\mathrm{D}_{\mathrm{KL}}}\big{(}q(\mathbf{x})\|p_{\theta}(\mathbf{x})\big{)}}_{\textrm{entropy + likelihood}}+\underbrace{{\mathrm{D}_{\mathrm{KL}}}\big{(}q_{\varphi}(\mathbf{z}_{0:1}|\mathbf{x})\|p_{\theta}(\mathbf{z}_{0:1}|\mathbf{x})\big{)}}_{\textrm{posterior error}}.$ (41)
+Thus, by minimizing the variational objective in diffusion models, we address two aspects: the likelihood of the data and the posterior error.
+From this perspective, it is quite a natural choice to also optimize the variational distribution $q_{\varphi}(\mathbf{z}_{0:1}|\mathbf{x})$ to match the model’s posterior distribution $p_{\theta}(\mathbf{z}_{0:1}|\mathbf{x})$.
+In other words, by making the forward process learnable, we optimize not only the reverse process to match the forward process but also allow the forward process to match the reverse process.
+This approach simplifies the task of matching the data distribution by the reverse process.
+B.2 Calculation of the Time Derivative
+As discussed in Section 3.1, the definition of forward process requires access to the time derivative of the function $F_{\varphi}$.
+This derivative is crucial for constructing a conditional ODE (see eq.
+8):
+$f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})=\left.\frac{\partial F_{\varphi}(\varepsilon,t,\mathbf{x})}{\partial t}\right|_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}.$ (42)
+First, let us clarify this notation: to find the ODE drift term $f_{\varphi}$ at point $\mathbf{z}_{t}$ and time $t$ conditionally on $\mathbf{x}$, we first apply the inverse function $F_{\varphi}^{-1}$ to identify $\varepsilon$, which corresponds to $\mathbf{z}_{t}$, $t$, and $\mathbf{x}$.
+Subsequently, we compute the partial derivative of $F_{\varphi}$ with respect to $t$ at $\varepsilon$ and $\mathbf{x}$.
+For certain parameterizations of $F_{\varphi}$, the time derivative can be determined analytically.
+For instance, $F_{\varphi}$ associated with conventional diffusion models (see Section E.1) have time derivatives available in a closed form.
+However, in the general case, the time derivative is not readily available.
+Fortunately, we can employ automatic differentiation tools like PyTorch [41] or JAX [5].
+The naive approach to determining the time derivative involves running a backpropagation procedure for each output of $F_{\varphi}$ and reconstructing each coordinate of the time derivative.
+Unfortunately, this method is not scalable for high-dimensional $F_{\varphi}$.
+Instead of backpropagation, which corresponds to a Vector Jacobian Product (VJP), we can use a forward differentiation mode, or Jacobian Vector Product (JVP).
+By fixing the inputs $\varepsilon$ and $\mathbf{x}$ of the function $F_{\varphi}$ and focusing solely on the input $t$, the Jacobian of $F_{\varphi}$ appears as a vector.
+Consequently, computing a JVP with a unit one-dimensional vector precisely produces the time derivative of $F_{\varphi}$.
+This approach allows for scalable computation of time derivatives.
+B.3 Calculation of the Conditional Score Function
+In Section 3.1, we define the forward process as a conditional SDE (eq. 9):
+$d\mathbf{z}_{t}$ $=\tilde{f}^{F}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w},\quad\textrm{where}$ (43)
+$\tilde{f}^{F}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})+\frac{g_{\varphi}^{2}(t)}{2}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x}).$
+To construct this SDE, access to the conditional score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is required.
+This section discusses methods for calculating this score function.
+First, consider the case where $F_{\varphi}$ is linear in $\varepsilon$ (further details on this parameterization are discussed in Section B.5):
+$F_{\varphi}(\varepsilon,t,\mathbf{x})=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t),$ (44)
+where $\mu_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ and $\sigma_{\varphi}:\mathbb{R}^{d}\times[0,1]\mapsto\mathbb{R}^{D}$, with the product $\varepsilon\sigma_{\varphi}(\mathbf{x},t)$ being element-wise.
+Consequently, the marginal probability densities of the forward process are Gaussian:
+$q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\mu_{\varphi}(\mathbf{x},t),\sigma_{\varphi}^{2}(\mathbf{x},t)).$ (45)
+In the case of Gaussian distributions, the score function can be easily found as:
+$\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\frac{\mu_{\varphi}(\mathbf{x},t)-\mathbf{x}}{\sigma_{\varphi}^{2}(\mathbf{x},t)}=\left.-\frac{\varepsilon}{\sigma_{\varphi}(\mathbf{x},t)}\right|_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}.$ (46)
+Now, consider the more general case where the distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is defined implicitly through an invertible transformation $F_{\varphi}$ of the random variable $\varepsilon\sim q(\varepsilon)$ into $\mathbf{z}_{t}$.
+To find the score function, we apply the change of variable formula:
+$\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\nabla_{\mathbf{z}_{t}}\left[\log q(\varepsilon)\Big{|}_{\varepsilon=F_{\varphi}^{-1}(\mathbf{z}_{t},t,\mathbf{x})}+\log\left|\frac{\partial F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})}{\partial\mathbf{z}_{t}}\right|\right].$ (47)
+The first part of the equation is the log-density of the noise distribution, which is straightforward to calculate.
+The second part is the log-determinant of the Jacobian matrix of the inverse transformation, $\log|J_{F}^{-1}|$.
+If we have access to $\log|J_{F}^{-1}|$, we can calculate $\log q\varphi(\mathbf{z}_{t}|\mathbf{x})$ using the change of variable formula and subsequently compute the gradient using automatic differentiation tools such as PyTorch [41] or JAX [5].
+The log-determinant $\log|J_{F}^{-1}|$ can be easily calculated when $F_{\varphi}$ is linear in $\varepsilon$, a case already considered above.
+For nonlinear, low-dimensional $F_{\varphi}$, the Jacobian $J_{F}^{-1}$ can be computed using $\mathcal{O}(D)$ backpropagations, and the log-determinant subsequently calculated.
+However, this approach is not scalable for high-dimensional cases.
+It is challenging to compute the log-determinant $\log|J_{F}^{-1}|$ in general high-dimensional cases, posing a limitation to the NFDM framework as it restricts the parameterization of the forward process to functions with accessible log-determinants.
+Nevertheless, the forward process can be parameterized with functions $F_{\varphi}$ that inherently provide access to the log-determinant $\log|J_{F}^{-1}|$ by design.
+For example, normalizing flow architectures [14, 28] facilitate the construction of forward processes with non-Gaussian, learnable distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
+B.4 Stochastic and Deterministic Sampling
+In Section 3.2, we define the reverse process as the following SDE (see eq. 11):
+$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}.$ (48)
+Let us now rewrite this equation, substituting the definition of $\tilde{f}^{B}_{\varphi}$ from eq. 10:
+$d\mathbf{z}_{t}=\left[f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\frac{g_{\varphi}^{2}(t)}{2}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\right]dt+g_{\varphi}(t)d\bar{\mathbf{w}}.$ (49)
+This expression represents the full form of the generative SDE, for which we derive the objective in Section A.1.
+After training, to sample data points from the models, we can simulate this SDE as outlined in eq.
+49 and depicted in Algorithm 2.
+However, as evident from eq.
+49, the generative process has a very simple dependence on $g_{\varphi}(t)$, which determines the level of stochasticity.
+Therefore, in practice, we can adjust the stochasticity level by modifying the function $g_{\varphi}(t)$.
+In the extreme case where $g_{\varphi}(t)\equiv 0$, the generative SDE completely loses its stochasticity and becomes an ODE:
+$d\mathbf{z}_{t}=f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}dt.$ (50)
+This ODE allows for an interpretation of the model as a Continuous Normalizing Flow (CNF) [6, 16], consequently enabling deterministic sampling from a pre-trained model and likelihood estimation.
+B.5 Parameterization of the Forward Process
+The parameterization of the variance function $g_{\varphi}(t)$ (eq.
+9) and the data point predictor $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ (eq.
+11) is straightforward.
+In our experiments, we parameterize them directly using neural networks.
+However, parameterizing the transformation $F_{\varphi}$ (eq.
+7) is more complex.
+NFDM requires $F_{\varphi}$ to be differentiable and invertible with respect to $\varepsilon$, as well as provide access to the log-determinant of the Jacobian matrix (see Section B.3).
+We propose two parameterizations for the forward process of NFDM.
+First, we propose a parameterization of $F_{\varphi}$ that is linear in $\varepsilon$:
+$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t),\quad\textrm{where}$ (51)
+$\mu_{\varphi}(\mathbf{x},t)$ $=(1-t)\mathbf{x}+t(1-t)\bar{\mu}_{\varphi}(\mathbf{x},t),$ (52)
+$\sigma_{\varphi}(\mathbf{x},t)$ $=\delta^{1-t}\left(\bar{\sigma}_{\varphi}(\mathbf{x},t)\right)^{t(1-t)}.$ (53)
+Here, $\bar{\mu}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ and $\bar{\sigma}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}_{+}$, with $\delta$ as a constant.
+This parameterization models $q(\mathbf{z}_{t}|\mathbf{x})$ as a conditional Gaussian distribution:
+$q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\mu_{\varphi}(\mathbf{x},t),\sigma_{\varphi}^{2}(\mathbf{x},t)).$ (54)
+In our experiments, we set $\delta^{2}=10^{-4}$.
+If $\bar{\mu}_{\varphi}$ is differentiable and $\bar{\sigma}_{\varphi}$ is differentiable and non-negative, this parameterization meets the requirements on $F_{\varphi}$.
+The functions $\mu_{\varphi}(\mathbf{x},t)$ and $\sigma_{\varphi}(\mathbf{x},t)$ are specifically designed such that at $t=0$, they equal $\mu_{\varphi}(\mathbf{x},0)=\mathbf{x}$ and $\sigma_{\varphi}(\mathbf{x},t)=\delta$, and at $t=1$, they become $\mu_{\varphi}(\mathbf{x},1)=0$ and $\sigma_{\varphi}(\mathbf{x},t)=I$.
+At intermediate times $10<t<$ 1, these functions can take arbitrary values, thereby ensuring that $q(\mathbf{z}_{0}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{0};\mathbf{x},\delta^{2}I)$ and $q(\mathbf{z}_{1}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{1};0,I)$, with arbitrary Gaussian distributions at intermediate steps.
+These restrictions are not necessary for $F_{\varphi}$, however, it eliminates the need to optimize the reconstruction loss ${\mathcal{L}_{\mathrm{rec}}}$ and the prior loss ${\mathcal{L}_{\mathrm{prior}}}$, as they become independent of the parameters $\varphi$ and $\theta$ (see Section A.1).
+Secondly, we propose a more general parameterization:
+$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=(1-t)\mathbf{x}+\big{(}\delta+(1-\delta)t\big{)}\bar{F}_{\varphi}(\varepsilon,t,(1-t)\mathbf{x}),$ (55)
+where $\bar{F}_{\varphi}:\mathbb{R}^{D}\times[0,1]\times\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$ is an invertible and differentiable function that allows access to the log-determinant of the Jacobian matrix.
+If $\bar{F}_{\varphi}$ is sufficiently flexible (e.g., a neural network with a normalizing flow architecture), it can parameterize non-Gaussian distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
+The additive term $(1-t)\mathbf{x}$ provides an inductive bias that shifts the mean of the distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ towards $\mathbf{x}$ when $t$ is near 00.
+The linear function $\delta+(1-\delta)t$ scales the distribution from a factor of $\delta$ at $t=0$ to $1$ at $t=1$, also imparting an inductive bias for increasing variance over time.
+The scaling $(1-t)\mathbf{x}$ in the third argument of $\bar{F}_{\varphi}$ ensures that $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ at $t=1$ does not depend on $\mathbf{x}$, allowing the reverse process to start from $p(\mathbf{z}_{1})=q_{\varphi}(\mathbf{z}_{1}|\mathbf{0})$ and thereby guaranteeing matching of prior distributions between the forward and reverse processes.
+Consequently, this eliminates the need for optimizing the prior loss ${\mathcal{L}_{\mathrm{prior}}}$.
+However, unlike the previous parameterization that guarantees $q(\mathbf{z}_{0}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{0};\mathbf{x},\delta^{2}I)$, in this case, there are no guarantees on the form of $q(\mathbf{z}_{0}|\mathbf{x})$, necessitating the optimization of the reconstruction loss ${\mathcal{L}_{\mathrm{rec}}}$ with respect to parameters $\varphi$ (see Section A.1).
+Importantly, even the Gaussian parameterization of $F_{\varphi}$ in eq.
+51 offers a significantly more flexible forward process compared to prior works.
+Although linear in $\varepsilon$, it may exhibit complex non-linear dependencies on $\mathbf{x}$ and $t$.
+Furthermore, we emphasize that the parameterizations of $F_{\varphi}$, $g_{\varphi}(t)$, and $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ discussed here are not the only possible ones and may be sub-optimal.
+For the purposes of this paper, we maintain a simple setup and leave the exploration of more effective parameterizations for future research.
+Appendix C Details of NFBM
+C.1 Parameterization of the Forward Process
+To parameterize the forward process of NFBM, we adopt the Gaussian parameterization discussed for NFDM in Section B.5:
+$F_{\varphi}(\varepsilon,t,\mathbf{x}_{0},\mathbf{x}_{1})$ $=\mu_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t)+\varepsilon\sigma_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t),\quad\textrm{where}$ (56)
+$\mu_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t)$ $=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}+t(1-t)\bar{\mu}_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t),$ (57)
+$\sigma_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t)$ $=\delta\left(\bar{\sigma}_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t)\right)^{t(1-t)}.$ (58)
+Here, $\bar{\mu}_{\varphi}:\mathbb{R}^{D}\times\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ and $\bar{\sigma}_{\varphi}:\mathbb{R}^{D}\times\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}_{+}$, with $\delta^{2}=10^{-4}$ in our experiments.
+Similar to the NFDM, this parameterization models $q(\mathbf{z}_{t}|\mathbf{x})$ as a conditional Gaussian distribution:
+$q_{\varphi}(\mathbf{z}_{t}|\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(\mathbf{z}_{t};\mu_{\varphi}(\mathbf{x}_{0},\mathbf{x}_{1},t),\sigma_{\varphi}^{2}(\mathbf{x}_{0},\mathbf{x}_{1},t)).$ (59)
+In contrast to the parameterization of NFDM, here $\mu_{\varphi}(\mathbf{x},t)$ and $\sigma_{\varphi}(\mathbf{x},t)$ are designed to ensure that $q(\mathbf{z}_{0}|\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(\mathbf{z}_{0};\mathbf{x}_{0},\delta^{2}I)$ and $q(\mathbf{z}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(\mathbf{z}_{1};\mathbf{x}_{1},\delta^{2}I)$.
+At intermediate times, $10<t<$ 1, $q(\mathbf{z}_{t}|\mathbf{x}_{0},\mathbf{x}_{1})$ may still be an arbitrary Gaussian distribution.
+Therefore, with this parameterization of NFBM, we do not need to optimize the reconstruction loss ${\mathcal{L}_{\mathrm{rec}}}$ and the prior loss ${\mathcal{L}_{\mathrm{prior}}}$, as they are independent of the parameters $\varphi$ and $\theta$ (see Section A.2).
+As with NFDM, the parameterization of NFBM that we provide in this section represents a simple design choice.
+We believe there are more efficient ways to parameterize the forward process of NFBM.
+However, in this paper, we want to focus on presenting the framework itself, rather than dwelling on design choices.
+C.2 Parameterization of the Reverse Process
+In Section 4, we define the reverse process of NFBM by the following SDE:
+$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}^{0,1}_{\theta}(\mathbf{z}_{t},t)\big{)}.$ (60)
+Here, the function $\hat{\mathbf{x}}^{0,1}_{\theta}(\mathbf{z}_{t},t)$ predicts both $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$.
+Assuming $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x}_{0},\mathbf{x}_{1})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})\approx p(\mathbf{z}_{1}|\mathbf{x}_{1})$, this formulation leads to the objective (see Section C.1):
+$\mathcal{L}=\mathbb{E}_{u(t)q(\mathbf{x}_{0},\mathbf{x}_{1})q_{\varphi}(\mathbf{z}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{1})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x}_{0},\mathbf{x}_{1})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right].$ (61)
+As previously mentioned, alternatively, we could define the reverse process as being fully conditional on the initial data point $\mathbf{x}_{1}$. This leads to the following SDE:
+$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}^{0}_{\theta}(\mathbf{z}_{t},t,\mathbf{x}_{1}),\mathbf{x}_{1}\big{)}.$ (62)
+This approach is akin to classifier-free guidance, where each step of the generative process depends on conditioning $\mathbf{x}_{1}$.
+Under the same assumptions—that $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x}_{0},\mathbf{x}_{1})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})\approx p(\mathbf{z}_{1}|\mathbf{x}_{1})$—this definition leads to a slightly different objective (the derivation is not provided here but can be easily inferred by following the steps described in Section A.2 for the previous definition):
+$\mathcal{L}=\mathbb{E}_{u(t)q(\mathbf{x}_{0},\mathbf{x}_{1})q_{\varphi}(\mathbf{z}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{1})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x}_{0},\mathbf{x}_{1})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t,\mathbf{x}_{1})\big{\|}_{2}^{2}\right].$ (63)
+Intuitively, the difference between these two definitions implies that the first reverse process, at a point $\mathbf{z}_{t}$ at time $t$, follows the average trajectory of forward processes passing through $\mathbf{z}_{t}$.
+Conversely, the second reverse process tracks the average trajectory of forward processes that pass through $\mathbf{z}_{t}$ and terminate at $\mathbf{z}_{1}$.
+The second approach may be beneficial when dealing with joint data distributions $q(\mathbf{x}_{0},\mathbf{x}_{1})$.
+However, in this paper, as we work with unpaired distributions $q(\mathbf{x}_{0})q(\mathbf{x}_{1})$, we have opted for the first approach in Section 4.
+Appendix D NFDM with restrictions
+D.1 Additional Discussion of NFDM with Restrictions
 In Section 5, we discussed training the NFDM with restrictions on the reverse process.
-This section aims to explore the limitations of such restrictions and to further discuss the training of NFDM with curvature penalties.
-When training NFDM with certain restrictions, it is crucial that these restrictions are feasible.
-For example, imposing overly stringent constraints on the reverse process (such as fixing its drift term $\hat{f}\equiv 0$) will render the forward process incapable of adapting, regardless of its flexibility.
-Similarly, unattainable penalties, like those on the squared norm of the drift term $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}$, will lead to biased solutions, as it’s impossible to match two processes when $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}\equiv 0$.
-Hence, in Section 4.1, we suggest penalizing the curvature of deterministic generative trajectories specifically with penalty ${\mathcal{L}_{\mathrm{crv}}}$ (eq.
-26).
-This is because for non-degenerate data distributions, there exist mappings that transform them into unit Gaussian noise along straight line trajectories.
+This section aims to further explore the training under restrictions and to examine the limitations of this technique.
+Typically, there is an infinite number of forward and reverse processes that correspond to each other.
+So how to choose between them?
+The standard NFDM, as described in Section 3, learns just one such pair of processes.
+However, the flexibility of NFDM opens up the possibility of choosing specific pairs of processes depending on the task.
+Suppose we aim to learn a reverse process with some specific properties.
+To achieve this, we must impose certain restrictions on the reverse process.
+One approach is to encode these restrictions directly in the parameterization of the reverse process.
+Alternatively, we can introduce penalties on the reverse process that would penalize deviations from the desired properties.
+In both cases, we expect the forward process to adapt and align with the reverse process.
+Consequently, the reverse process will match the forward process and, as a result, the data distribution, while possessing the desired properties.
+In Section 5, we discussed the restriction of the reverse process to have straight deterministic generative trajectories.
+However, we may apply the same strategy to ensure arbitrary properties.
+We believe that the NFDM framework may help construct various new generative dynamics for a variety of tasks and domains.
+Nevertheless, training the NFDM with restrictions has limitations.
+The primary concern is the feasibility of these restrictions.
+For instance, imposing overly stringent constraints on the reverse process (such as fixing its drift term $\hat{f}\equiv 0$) will render the forward process incapable of adapting, regardless of its flexibility.
+Similarly, unattainable penalties, such as those on the squared norm of the drift term $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}$, will lead to biased solutions, as it is impossible to match two processes when $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}\equiv 0$.
+Additionally, even when a restriction is feasible, if the model parameterization is too simplistic, it may not be able to match processes and satisfy the restriction, leading to bias.
+However, in our experiments, we find that even the Gaussian parameterization of the forward process (see Section B.5) is flexible enough in practice to successfully learn various restrictions, such as straight generative trajectories.
+D.2 Details of NFDM with Curvature Penalty
+In Section 5, we suggest penalizing the curvature of deterministic generative trajectories specifically with the ${\mathcal{L}_{\mathrm{crv}}}$ (eq. 16) penalty.
 To calculate the curvature penalty ${\mathcal{L}_{\mathrm{crv}}}$ (eq. 16), we proceed as follows. First, we use the chain rule to rewrite the time derivative:
@@
 65 is the time derivative of a function.
-As discussed in section 3.1, the time derivatives can be determined as a Jacobian Vector Product (JVP)with a one-dimensional unit vector.
+As discussed in section B.2, the time derivatives can be determined as a Jacobian Vector Product (JVP)with a one-dimensional unit vector.
 Notably, the first term in eq.
@@
 $v=\begin{bmatrix}\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\\ 1\end{bmatrix}.$ (66)
-Consequently, ${\mathcal{L}_{\mathrm{crv}}}$ can be computed as the JVP of $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)$ with the vector $v$.
-Appendix B Connections of NFDM with other approaches
-B.1 Special cases of NFDM
+Consequently, ${\mathcal{L}_{\mathrm{crv}}}$ can be efficiently computed as the JVP of $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)$ with the vector $v$.
+Additionally, we would like to note a couple of aspects of learning the NFDM with a curvature penalty.
+First, unlike the regular NFDM, the curvature penalty necessitates that $F_{\varphi}$ be twice differentiable with respect to time.
+Second, although we parameterize $\hat{f}_{\theta,\varphi}$ (eq.
+11) through $\tilde{f}^{B}_{\varphi}$ (eq.
+10), ${\mathcal{L}_{\mathrm{crv}}}$ influences the forward process but does not penalize the curvature of the conditional forward trajectories (see illustrations in Section G.1).
+Appendix E Connections of NFDM with Other Approaches
+In this section we continue discussing connections between NFDM and related works.
+E.1 Special Cases of NFDM
 Many existing works define the forward process in diffusion models as either fixed or simply parameterized processes.
@@
 Soft Diffusion.
-Daras et al.
-(2022) consider a more general case of fixed forward processes, $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};C_{t}\mathbf{x},\sigma_{t}I)$, which can be parameterized as $F(\varepsilon,t,\mathbf{x})=C_{t}\mathbf{x}+\sigma_{t}\varepsilon$.
+[9] consider the case of a fixed forward processes $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};C_{t}\mathbf{x},\sigma_{t}I)$, which can be reparameterized as $F(\varepsilon,t,\mathbf{x})=C_{t}\mathbf{x}+\sigma_{t}\varepsilon$.
 Such distributions include, for example, combinations of blurring and the injection of Gaussian noise.
 Star-Shaped Diffusion.
-Okhotin et al.
-(2023) extended conventional diffusion models to include distributions from the exponential family.
+[40] extended conventional diffusion models to include distributions from the exponential family.
 Although Star-Shaped Diffusion is a discrete-time approach and does not directly correspond to NFDM, the latter can work with exponential family distributions through reparameterization functions.
@@
 Variational Diffusion Models.
-Kingma et al.
-(2021) proposed forward conditional distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ as $\mathcal{N}(\mathbf{z}_{t};\alpha_{\varphi}(t)\mathbf{x},\sigma_{\varphi}^{2}(t)I)$ with learnable parameters $\varphi$.
-In the context of NFDM, this distribution can be parameterized by $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{\varphi}(t)\mathbf{x}+\sigma_{\varphi}(t)\varepsilon$.
+[27] proposed forward conditional distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ as $\mathcal{N}(\mathbf{z}_{t};\alpha_{\varphi}(t)\mathbf{x},\sigma_{\varphi}^{2}(t)I)$ with learnable parameters $\varphi$.
+In the context of NFDM, this can be realized by $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{\varphi}(t)\mathbf{x}+\sigma_{\varphi}(t)\varepsilon$.
 LSGM.
-Vahdat et al.
-(2021) suggest an alternative approach for parameterizing the forward process, proposing diffusion in a latent space of a VAE.
+[65] suggest an alternative approach for parameterizing the forward process, proposing diffusion in a latent space of a VAE.
 Therefore, the forward process is characterized by a distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}E_{\varphi}(\mathbf{x}),\sigma_{t}^{2}I)$, where $E_{\varphi}$ is the encoder of the VAE.
@@
 To align the reverse process, the reconstruction distribution should be $p(\mathbf{x}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{x};D_{\varphi}(\mathbf{z}_{0}),\delta^{2}I)$, where $D_{\varphi}$ is VAE’s decoder.
-DiffEnc.
-Nielsen et al.
-(2023) proposed a more general forward process, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}f_{\varphi}(\mathbf{x},t),\sigma_{t}^{2}I)$, which, unlike LSGM, transforms $\mathbf{x}$ in a time-dependent manner.
-This forward process can also be described in terms of NFDM as $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}f_{\varphi}(\mathbf{x},t)+\sigma_{t}\varepsilon$.
-B.2 Connections of NFDM objective
-In this section, we delve into the details of the NFDM’s objective function (eq. 18), with a particular focus on the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ (eq. 20).
-First, let’s unpack the ${\mathcal{L}_{\mathrm{diff}}}$ by substituting the definitions of $\tilde{f}^{B}_{\varphi}$ and $\hat{f}_{\theta,\varphi}$:
-${\mathcal{L}_{\mathrm{diff}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\right\|_{2}^{2}\right]$ (43)
-$=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\underbrace{\Big{(}f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\Big{)}}_{\textrm{Flow Matching term}}+\frac{g_{\varphi}^{2}(t)}{2}\underbrace{\Big{(}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\Big{)}}_{\textrm{Score Matching term}}\right\|_{2}^{2}\right].$ (44)
-This formulation clearly delineates two components of the objective: the first calculates the difference between the ODE drift terms, and the second calculates the difference between the score functions.
-Moreover, this expression highlights the role of $g_{\varphi}(t)$.
-When $g_{\varphi}(t)$ is small, the forward and reverse processes exhibit smoother trajectories, and the objective is dominated by the first term.
-Conversely, when $g_{\varphi}(t)$ is large, the processes exhibit more stochastic trajectories, and the objective is dominated by the second term.
-Crucially, in the extreme scenarios where $g_{\varphi}(t)$ approaches either 00 or $\infty$, the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ corresponds to either a reweighted Flow Matching loss (Lipman et al., 2022) or a reweighted Score Matching loss (Vincent, 2011), respectively:
-$\lim_{g_{\varphi}(t)\rightarrow 0}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow 0}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\right\|_{2}^{2}\right],$ (45)
-$\lim_{g_{\varphi}(t)\rightarrow\infty}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow\infty}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{g_{\varphi}^{2}(t)}{4}\left\|\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\right\|_{2}^{2}\right]$ (46)
-B.3 Connections of NFDM with Stochastic Interpolants
-In this section, we explore the connections between NFDM and Stochastic Interpolants, as proposed by Albergo et al.
-(2023).
+NDM and DiffEnc.
+[4, 39] proposed a more general forward process, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}f_{\varphi}(\mathbf{x},t),\sigma_{t}^{2}I)$, which, unlike LSGM, transforms $\mathbf{x}$ in a time-dependent manner.
+This forward process can also be described in terms of NFDM as the special case $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}f_{\varphi}(\mathbf{x},t)+\sigma_{t}\varepsilon$.
+Curvature minimization.
+[31] suggested learning the distribution $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ of the forward process at time step $t=1$ as a conditional Gaussian.
+They parameterize the forward process via linear interpolation between $\mathbf{x}$ and $\mathbf{z}_{1}$.
+Within the NFDM framework, we can express this as $F_{\varphi}(\varepsilon,t,\mathbf{x})=(1-t)\mathbf{x}+t\big{(}\mu_{\varphi}(\mathbf{x})+\sigma_{\varphi}(\mathbf{x})\varepsilon\big{)}$.
+Subsequently, the authors train their model by optimizing the Flow Matching objective in conjunction with minimizing the Kullback–Leibler (KL) divergence between $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ and $p(\mathbf{z})$.
+This objective corresponds to minimization of diffusion term ${\mathcal{L}_{\mathrm{diff}}}$ and prior term ${\mathcal{L}_{\mathrm{prior}}}$ of NFDM objective (see Section A.1) for the forward process proposed by [31].
+E.2 Compatible and Orthogonal Approaches
+Despite the simulation-free nature of the training procedure, diffusion models still necessitate full reverse process simulations for sample generation, leading to slow and computationally expensive inference.
+To address this, in an orthogonal line of works alternative sampling methods have been studied, such as deterministic sampling in discrete [56] and continuous time [60] and novel numerical solvers [61, 34, 37, 52].
+Additionally, distillation techniques have been applied to both discrete [49] and continuous time models [57, 73, 71] to enhance sampling speed, albeit at the cost of training additional models.
+Most of these approaches are compatible with NFDM and may be adapted for further gains.
+The significance of deterministic trajectories in efficient sampling was highlighted by [23].
+At the same time, building on diffusion model concepts, [32, 36, 2] introduced simulation-free methods for learning deterministic generative dynamics.
+Based on these ideas, [36] and [35] proposed distillation procedures to straighten deterministic generative trajectories.
+Since these methods are based on distillation, we consider them orthogonal to NFDM.
+E.3 Other Related Works
+Schrödinger Bridge models.
+Recent studies [72, 10, 53] have explored generative models based on Schrödinger Bridge theory and finite-time diffusion constructions.
+Unlike NFDM, these models deviate from the simulation-free paradigm by necessitating the full simulation of stochastic processes for training, which significantly increases their computational cost.
+Moreover, these studies typically aim to address the optimal stochastic control problem, which is outside the scope of our work.
+Nevertheless, it is important to highlight that within this research domain, some authors have proposed advanced techniques for parameterizing learnable conditional processes.
+For instance, in [33], the authors introduced the use of splines to parameterize the conditional process, thereby constructing a more flexible and efficient model.
+This parameterization technique can be seen as a specific instantiation of the parameterization procedure proposed by the NFDM framework.
+Minibatch optimal transport.
+[43] and [63] proposed to construct the forward process with optimal data-noise couplings to learn straighter generative trajectories.
+While this method is theoretically justified, it relies on estimating an optimal coupling over minibatches of the entire dataset, which, for large datasets, may become uninformative as to the true optimal transport coupling.
+In contrast, NFDM-OT directly penalizes curvature of the generative trajectories.
+Stochastic Interpolants.
+In this section, we explore the connections between NFDM and Stochastic Interpolants, as proposed by [1].
 Both NFDM and Stochastic Interpolants introduce a more general family of forward processes through the use of a reparameterization function.
@@
 Furthermore, these objectives are tailored towards learning the velocity field and the score function, rather than optimizing likelihood, as is the case with NFDM.
-In their practical applications, Stochastic Interpolants are demonstrated with only simple parameterizations of the forward process.
+In their practical applications, Stochastic Interpolants are demonstrated with a simple parameterizations of the forward process.
 They propose a theoretical framework for learning the forward process that would result in an reverse process characterized by dynamical optimal transport.
@@
 Moreover, their work does not clearly articulate how Stochastic Interpolants might be applied to learning other generative dynamics.
-In contrast, NFDM introduces a more generalized method for learning generative dynamics.
+In contrast, NFDM introduces a more generic method for learning generative dynamics.
 Moreover, when NFDM is learned with restrictions or penalties (see Section 5), it remains within a min-min optimization paradigm, which can be addressed more efficiently.
-Appendix C Implementation details
-Our evaluation of NFDM includes tests on synthetic data, CIFAR-10 (Krizhevsky et al., 2009), and two downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) datasets.
-To maintain consistency with baselines, we employ horizontal flipping as a data augmentation technique in training models on CIFAR-10 and ImageNet (Song et al., 2020b, 2021).
-For density estimation of discrete data, uniform dequantization is used (see Section C.1).
+E.4 Connections of NFDM Objective
+In this section, we delve into the details of the NFDM’s objective function (eq.
+12).
+Specifically we will consider the diffusion term of the objective ${\mathcal{L}_{\mathrm{diff}}}$ (see Section A.1).
+First, let’s unpack the ${\mathcal{L}_{\mathrm{diff}}}$ by substituting the definitions of $\tilde{f}^{B}_{\varphi}$ and $\hat{f}_{\theta,\varphi}$:
+${\mathcal{L}_{\mathrm{diff}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}$ $\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\right\|_{2}^{2}\right]$ (67)
+$=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}$ $\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\underbrace{\Big{(}f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\Big{)}}_{\textrm{Flow Matching term}}+\right.\right.$ (68)
+$\quad\left.\left.\frac{g_{\varphi}^{2}(t)}{2}\underbrace{\Big{(}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\Big{)}}_{\textrm{Score Matching term}}\right\|_{2}^{2}\right].$
+This formulation clearly delineates two components of the objective: the first calculates the difference between the ODE drift terms, and the second calculates the difference between the score functions.
+Moreover, this expression highlights the role of $g_{\varphi}(t)$.
+When $g_{\varphi}(t)$ is small, the forward and reverse processes exhibit smoother trajectories, and the objective is dominated by the first term.
+Conversely, when $g_{\varphi}(t)$ is large, the processes exhibit more stochastic trajectories, and the objective is dominated by the second term.
+Crucially, in the extreme scenarios where $g_{\varphi}(t)$ approaches either 00 or $\infty$, the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ corresponds to either a reweighted Flow Matching loss [32] or a reweighted Score Matching loss [67], respectively:
+$\lim_{g_{\varphi}(t)\rightarrow 0}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow 0}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\right\|_{2}^{2}\right],$ (69)
+$\lim_{g_{\varphi}(t)\rightarrow\infty}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow\infty}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{g_{\varphi}^{2}(t)}{8}\left\|\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\right\|_{2}^{2}\right]$ (70)
+This connection highlights the importance of optimizing the full objective in eq.
+67, which contains both Flow Matching and Score Matching components.
+Simplifying the objective to include only one component renders it intractable from a variational perspective.
+Specifically, when $g_{\phi}(t)$ approaches either 00 or $\infty$, the objective, a variational bound on likelihood, becomes infinitely large.
+There are two primary reasons why many popular diffusion models are trained using just one of these components.
+Firstly, many diffusion models employ a simple linear parameterization of the forward process.
+For these models, both the conditional vector field $f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ and the conditional score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ are merely linear combinations of $\mathbf{z}_{t}$ and $\mathbf{x}$.
+As [26] demonstrated, for such models, Flow Matching and Score Matching can be viewed as reweighted ELBO.
+However, for more general non-linear forward process in NFDM, optimizing Flow Matching and Score Matching objectives is not equivalent.
+Secondly, in other models, the Flow Matching and Score Matching objectives are derived under the assumption that the forward process is fixed.
+Without this assumption, we cannot derive either objective in the general case.
+In NFDM, we learn the forward process end-to-end with the reverse process.
+Thus, reducing the NFDM’s objective could lead to the model’s collapse, as it is not the correct objective to optimize.
+We would also like to emphasize that the NFDM framework, including its objective in eq.
+67, represents a generalization of conventional diffusion models, rather than an alternative approach to training.
+Consequently, if we take an existing forward process from conventional diffusion models and incorporate it into the objective $\mathcal{L}$, we will not create a new model or optimization procedure.
+Instead, this will align with approaches that focus on likelihood-based training of conventional diffusion models, as explored in [58, 74].
+NFDM facilitates the convenient pre-specification of fixed forward processes.
+Hence, it might be utilized to define a fixed forward process and subsequently train it with the Flow Matching objective.
+However, this is a topic for future research and is beyond the scope of this work.
+Appendix F Implementation Details
+Table 4: Training hyper-parameters.
+CIFAR-10 ImageNet 32 ImageNet 64 AFHQ 64
+Channels 256 256 192 192
+Depth 2 3 3 3
+Channels multipliers 1,2,2,2 1,2,2,2 1,2,3,4 1,2,2,2
+Heads 4 4 4 4
+Heads Channels 64 64 64 64
+Attention resolution 16 16,8 32,16,8 32,16,8
+Dropout 0.0 0.0 0.0 0.0
+Effective Batch size 256 1024 2048 1024
+GPUs 2 4 16 16
+Epochs 1000 200 250 100
+Iterations 391k 250k 157k 102k
+Learning Rate 4e-4 1e-4 1e-4 1e-4
+Learning Rate Scheduler Polynomial Polynomial Constant Constant
+Warmup Steps 45k 20k - -
+Our evaluation of NFDM and NFBM includes tests on synthetic data, CIFAR-10 [30], two downsampled ImageNet [11, 66] datasets, and downsampled AFHQ [8].
+To maintain consistency with baselines, we employ horizontal flipping as a data augmentation technique in training models on CIFAR-10 and ImageNet [60, 58].
+For density estimation of discrete data, uniform dequantization is used (see Section F.1).
 We parameterize $\hat{\mathbf{x}}_{\theta}$ (eq.
-17) in the reverse process using a 5-layer perceptrons with 512 neurons in each layer for synthetic data and the U-Net architecture from Dhariwal & Nichol (2021) for images.
+11) in the reverse process using a 5-layer perceptrons with 512 neurons in each layer for synthetic data and the U-Net architecture from [12] for images.
 In all experiments, a 3-layer perceptrons with 64 neurons in each layer is employed to parameterize $g_{\varphi}$ (eq.
-12), and for $F_{\varphi}$ (eq.
-9), we use an identical neural network to that of $\hat{\mathbf{x}}_{\theta}$.
-The sole difference is that for $F_{\varphi}$, we double the output of the last layer to parameterize $\bar{\mu}_{\varphi}$ (eq.
-23) and $\bar{\sigma}_{\varphi}$ (eq.
-24) with same model (see Section 3.5).
+9).
+For the parameterization of $F_{\varphi}$(eq.
+7) in NFDM as Gaussian, we employ a neural network identical to that used for $\hat{\mathbf{x}}_{\theta}$.
+The only difference is that for $F_{\varphi}$, we double the output of the last layer to parameterize both $\bar{\mu}_{\varphi}$ (eq.
+52) and $\bar{\sigma}_{\varphi}$(eq.
+53) using the same model (see Section B.5).
+For parameterizing $F_{\varphi}$ (eq.
+13) in NFDM, we adopt the same approach but double the input size of the first layer to condition the forward process on both $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ (see Section C.1).
+To parameterize $F_{\varphi}$ (eq.
+7) in NFDM as non-Gaussian, we apply the Glow [28] architecture tailored to the corresponding dataset to transform $\varepsilon$ into $\mathbf{z}_{t}$.
+To incorporate conditioning, we use a neural network identical to that of $\hat{\mathbf{x}}_{\theta}$ to generate an embedding $u(\mathbf{x},t)$.
+We then add this conditioning to the input of each non-linear transformation in the coupling layers.
+To ensure the constraints $g_{\varphi}\geq 0$ (eq. 9) and $\bar{\sigma}_{\varphi}\geq 0$ (eqs. 53 and 58) are met, we apply the softplus function to the outputs of the neural networks.
 The models were trained using the Adam optimizer with the following parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, a weight decay of $0.0$, and $\epsilon=10^{-8}$.
@@
 The specific hyperparameters are detailed in Table 4.
-Training was carried out on Tesla V100 GPUs
-Table 3: Training hyper-parameters.
-CIFAR-10 ImageNet 32 ImageNet 64
-Channels 256 256 192
-Depth 2 3 3
-Channels multipliers 1,2,2,2 1,2,2,2 1,2,3,4
-Heads 4 4 4
-Heads Channels 64 64 64
-Attention resolution 16 16,8 32,16,8
-Dropout 0.0 0.0 0.0
-Effective Batch size 256 1024 2048
-GPUs 2 4 16
-Epochs 1000 200 250
-Iterations 391k 250k 157k
-Learning Rate 4e-4 1e-4 1e-4
-Learning Rate Scheduler Polynomial Polynomial Constant
-Warmup Steps 45k 20k -
-C.1 Dequantization
+Training was carried out on Tesla V100 GPUs.
+F.1 Dequantization
 For reporting the NLL, we employ standard uniform dequantization. The NLL is estimated using an importance-weighted average, given by
@@
 where $\mathbf{x}\in[0,\dots,255]$.
-C.2 Parameterization
-We define the parameterization of the function $F_{\varphi}$ as follows in eq. 22:
-$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t).$ (48)
-In addition to inferring $F_{\varphi}$, we require access to its inverse $F^{-1}_{\varphi}$ and the logarithm of the determinant of the inverse transformation $\log|J_{F}^{-1}|$ (eq.
-14).
-The calculation of the inverse is straightforward:
-$F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=\frac{\mathbf{z}_{t}-\mu_{\varphi}(\mathbf{x},t)}{\sigma_{\varphi}(\mathbf{x},t)}.$ (49)
-The logarithm of the determinant of the inverse transformation, $\log|J_{F}^{-1}|$, can also be readily determined:
-$\log|J_{F}^{-1}|=-\log|J_{F}|=-\sum_{i=1}^{D}\log\left[\sigma_{\varphi}(\mathbf{x},t)\right]_{i},$ (50)
-where $D$ represents the dimensionality of the data, and $\left[\sigma_{\varphi}(\mathbf{x},t)\right]_{i}$ refers to $i$-th output of the $\sigma_{\varphi}$ function.
-To ensure the constraints $g_{\varphi}\geq 0$ (eq. 12) and $\bar{\sigma}_{\varphi}\geq 0$ (eq. 24) are met, we apply the softplus function to the outputs of the neural networks.
+F.2 NFBM with Obstacles
+In Section 6.3, we demonstrate how the concept of learning NFDM with the restrictions discussed in Section 5 can be adapted to NFBM for learning generative dynamics that avoid obstacles.
+For this experiment, we propose training the NFBM models with an additional penalty that has higher values at points closer to obstacles.
+Specifically, we applied the following objective:
+$\mathcal{L}_{\textrm{obs}}=\mathbb{E}_{u(t)q(\mathbf{z}_{t})}[f(\mathbf{z}_{t})],$ (72)
+where $f(\mathbf{z}_{t})$ is a penalty function.
+In our experiments, we use the probability density of a mixture of Gaussian distributions, truncated at some low values to 00.
+This truncation is necessary to ensure that the model does not push trajectories infinitely away from the obstacle.
+In contrast to ${\mathcal{L}_{\mathrm{crv}}}$ proposed for NFDM-OT, which penalizes the reverse process influencing the forward process as a result, $\mathcal{L}_{\textrm{obs}}$ directly penalizes the forward process, thereby influencing the properties of the reverse process.
+Appendix G Additional Results
+G.1 Visualisations of Forward Trajectories of NFDM-OT
+Figure 3: First coordinates of forward deterministic trajectories of NFDM-OT started from points $\mathbf{x}=(-1,-1)$ and $\mathbf{x}=(1,1)$.
+In this section, we provide visualizations of the forward trajectories of NFDM-OT, as illustrated in the two-dimensional data distribution experiment shown in Figure 1(b).
+Figure 3 depicts the deterministic trajectories of NFDM-OT.
+It is easy to see that the forward process has learned some highly curved trajectories.
+This example highlights several key aspects of NFDM-OT and NFDM more broadly.
+Firstly, this example underscores the point discussed in Section D.2: while NFDM-OT penalizes the curvature of the reverse process, which is parameterized through the forward process (see Section 3.2), it does not penalize the curvature of the forward process itself.
+Secondly, this example demonstrates the importance of learning nonlinear forward processes.
+In attempting to adapt to the straight-line reverse process, the forward process ends up being highly curved, which in turn aids the reverse process.
+This emphasizes the importance of the NFDM framework, which facilitates the learning of such nonlinear forward processes that may help to learn better generative dynamics.
+Finally, we note that we parameterize the forward process with a Gaussian distributions featuring learnable mean and variance (see Section B.5).
+Even this relatively simple parameterization significantly enhances the flexibility of diffusion models compared to conventional diffusion approaches, enabling the learning of complex forward processes.
+G.2 Additional Results of NFDM and NFDM-OT on Image Datasets
+Table 5: Comparison of NFDM with NFDM-OT results on density estimation tasks. We present results in terms of BPD, lower is better.
+Model CIFAR10 ImageNet 32
+NFDM $2.49$ $3.36$
+NFDM-OT $\mathbf{2.62}$ $\mathbf{3.45}$
+Table 6: Comparison of NFDM with NFDM-OT results on few-step generation. We present results in terms of FID, lower is better.
+CIFAR-10 ImageNet 32
+Model NFE $\downarrow$ FID $\downarrow$ NFE $\downarrow$ FID $\downarrow$
+$4$ $50.12$ $4$ $57.60$
+NFDM $12$ $21.88$ $12$ $24.74$
+$4$ $7.76$ $4$ $6.13$
+NFDM-OT $12$ $5.20$ $12$ $4.11$
+In this section we provide additional results, that directly compare NFDM and NFDM-OT.
+This results extend Tables 1 and 2.
+As observed, NFDM without any additional restrictions provides a better log-likelihood estimation compared to NFDM-OT.
+However, as expected, NFDM exhibits much worse sample quality compared to NFDM-OT in a few steps generation.
+We attribute this property to the higher curvature of NFDM’s generative trajectories.
+G.3 Generated Samples
+In this section, we present generated samples from the NFDM, NFDM-OT, and NFBM models, trained on various datasets.
+NFDM: Figure 4 shows samples from the NFDM model with Gaussian $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ (see Section B.5).
+We provide samples from NFDM trained on the CIFAR-10, ImageNet 32, and ImageNet 64 datasets.
+NFDM-OT. Figure 5 displays samples from NFDM trained on ImageNet 64 dataset. We generate these samples by simulating the ODE across $12$, $4$, and $2$ steps.
+NFBM.
+Figure 6 demonstrates the generative trajectories of the NFBM model trained on the AFHQ 64 dataset.
+This model starts from an image of a dog at time step $t=1$ and simulate the SDE to time step $t=0$.
+(a) CIFAR-10
+(b) ImageNet 32
+(c) ImageNet 64
+Figure 4: Generated samples from NFDM trained on various datasets.
+[…]
+Figure 6: Generative trajectories from NFBM trained on AFHQ 64.

### v3

@@
 We leave it for future research to explore different parameterizations and modified objectives for NFDM that may yield even better results.
-Please refer to section G.3 for generated samples.
+Please refer to section G.3 for generated samples. The code is available at https://github.com/GrigoryBartosh/neural_diffusion.
 6.1 Likelihood Estimation
@@
 Nevertheless, conventional diffusion models, typically relying on a linear Gaussian forward process, may not optimally fit some data distributions.
-To address this, alternative forward processes have been explored, such as combining blurring with Gaussian noise injection [45, 9, 22], diffusion in the wavelet spectrum [42], and forward processes based on the exponential family [40].
+To address this, alternative forward processes have been explored, such as alternative linear transformations [16, 55], combining blurring with Gaussian noise injection [46, 9, 23], diffusion in the wavelet spectrum [43], and forward processes based on the exponential family [41].
 These models are limited by their fixed forward processes and may be seen as specific instances of NFDM.
@@
 NFDM outperforms baseline diffusion models on standard benchmarks, showcasing its effectiveness.
-However, NFDM is not merely just another model, but rather it is a versatile framework that facilitates the predefined specification and learning of a forward process.
+However, NFDM is not just another model, but rather it is a versatile framework that facilitates the predefined specification and learning of a forward process.
 Similarly, the penalty on the curvature of deterministic trajectories discussed in Section 5 represents just one specific example of the restrictions that NFDM can accommodate.
@@
 Once the forward process is parameterized using a neural network, this leads to approximately 2.2 times longer optimization iteration of NFDM compared to conventional diffusion models.
-Additionally, as discussed in Section 3.1, NFDM doesn’t allow parameterisation of the forward process with functions, that doesen’t match certain conditions.
+Additionally, as discussed in Section 3.1, NFDM is limited to forward process parameterisations with functions that match specific conditions.
 Despite these challenges, we are optimistic about the potential of NFDM.
@@
 To align the reverse process, the reconstruction distribution should be $p(\mathbf{x}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{x};D_{\varphi}(\mathbf{z}_{0}),\delta^{2}I)$, where $D_{\varphi}$ is VAE’s decoder.
+ShiftDDPM.
+[74] proposed learning a function that shifts the mean and covariance of $q_{\phi}(\mathbf{z}_{t}|\mathbf{x})$ by conditional information.
+In our work, we do not focus on conditional generation.
+However, we believe that the NFDM framework can be readily adapted for conditional generation by incorporating the conditioning variable $\mathbf{c}$ into both the transformation $F_{\phi}(\varepsilon,t,\mathbf{x},\mathbf{c})$ and the predictor $\hat{\mathbf{x}}(\mathbf{z}_{t},t,\mathbf{c})$.
 NDM and DiffEnc.
@@
 Figure 6: Generative trajectories from NFBM trained on AFHQ 64.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Our paper follows the structure of the abstract and introduction directly.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss limitations in general in Section 8 and in Section B.3 in more detail.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provide proofs in Appendix A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We provide detailed description of the model in Sections 3 and C.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: No new data has been generated, and the code will be made available late.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The data configuration follows from previous work, while most of the experimental details are provided in Appendix F.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: Computation of error bars is vary computationally demanding. We follow standard practice in the literature and report the average result.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide compute resources details in Appendix F.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We read and checked the code of ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: We don’t believe this work has any societal impact.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper does not pose such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All data and software used is properly credited and licenses are respected.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: No assets are introduced in the paper.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper did not require crowdsourcing.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No study participants were involved.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.